### v1

Abstract
Bayesian filtering is a key tool in many problems that involve the online processing of data, including data assimilation, optimal control, nonlinear tracking and others. Unfortunately, the implementation of filters for nonlinear, possibly high-dimensional, dynamical systems is far from straightforward, as computational methods have to meet a delicate trade-off involving stability, accuracy and computational cost. In this paper we investigate the design, and theoretical features, of constrained Bayesian filters for state space models. The constraint on the filter is given by a sequence of compact subsets of the state space that determines the sources and targets of the Markov transition kernels in the dynamical model. Subject to such constraints, we provide sufficient conditions for filter stability and approximation error rates with respect to the original (unconstrained) Bayesian filter. Then, we look specifically into the implementation of constrained filters in a continuous-discrete setting where the state of the system is a continuous-time stochastic Itô process but data are collected sequentially over a time grid. We propose an implementation of the constraint that relies on a data-driven modification of the drift of the Itô process using barrier functions, and discuss the relation of this scheme with methods based on the Doob $h$-transform. Finally, we illustrate the theoretical results and the performance of the proposed methods in computer experiments for a partially-observed stochastic Lorenz 96 model.
keywords:
Bayesian filters; filter stability; approximation errors; continuous-discrete filtering; Doob $h$-transform; constrained SDEs.
{MSCcodes}
62M20, 60G35, 60H10, 65C05
1 Introduction
1.1 Motivation and background
Bayesian filtering is a key tool in many problems that involve the online processing of data, including data assimilation [Leeuwen09, Law15, Reich15, Fearnhead18], optimal control [Bensoussan92, Mitter02, Georgiou13], nonlinear tracking [Arulampalam02b, Ristic04, Stone13], etc. The filtering problem is to compute the probability distribution of the state of a dynamical system conditional on available observations. The latter are collected over time, hence the distribution of the state has to be updated when new data becomes available [Anderson79, Bain08].
The exact implementation of Bayesian filters is rarely feasible beyond the Kalman filter for linear and Gaussian models [Kalman60, Kalman61, Anderson79, Ristic04]. In general, approximations are needed, which include Kalman-based methods [Anderson79, Evensen03, Julier04], recursive Monte Carlo algorithms [Gordon93, Doucet00, Evensen03, Djuric03], variational schemes [Smidl08, Frazier23] and combinations of them. All these methods suffer from limitations which are related to a complex trade-off involving stability, accuracy and computational cost. For example, Kalman-based methods become very inaccurate when the system of interest displays strong nonlinearities or is subject to heavy-tailed perturbations [Ristic04], and particle filters (PFs) [Gordon93, Kitagawa96, Doucet00, Djuric03] usually struggle when the observations are highly-informative or the dimension of the model dimension is high [Bengtsson08, Snyder08].
One approach to the implementation of robust-yet-accurate filters is the imposition of constraints on the dynamics of the system of interest. The key idea underlying this strategy is to restrict the exploration of the state space to typical solutions, while excluding others which are feasible but very unlikely. In this way, one expects to obtain algorithms which are numerically stable and yield good accuracy with high probability. In a broad sense, auxiliary [Pitt01, Douc09b, Branchini21] and twisted [Whiteley14, Ala16] PFs can be interpreted as constrained algorithms, as they adapt the sampling kernel using the likelihood (auxiliary PFs) or optimize it to minimize the variance of certain estimators (twisted PFs). Another way to constrain the filter dynamics is to truncate the support of the system state, a solution that has been explored methodologically for Kalman filters [Garcia12, Amor17] and also exploited to analyze filter stability [Crisan20] and the convergence of PFs [Heine08].
In this work, we are specifically interested in continuous-discrete filtering problems, where the state of the system of interest is a continuous-time stochastic process $\mathtt{X}(t)$ characterized by an Itô stochastic differential equation (SDE), and the observations $Y_{1},Y_{2},\ldots,Y_{n},$... are collected sequentially at specific time instants $t_{1},t_{2},\ldots,t_{n},$... [Akyildiz24]. In this setting, any constraint on the system dynamics needs to be translated into a modification of the underlying SDE. Such modifications are not straightforward. Related methods include guided SDEs, where an additional drift term is constructed using Doob’s $h$-transform [Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25]. This transform, denoted $h(x,t)$, is generally intractable, however, and it is common to resort to linearizations of the original SDE in order to approximate $h(x,t)$. The methods in [Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25] follow this approach. Moreover, they are developed for infinite-dimensional models and they enable sampling from regions of high probability given the observed data. The authors of [Chopin23], on the other hand, avoid linearizations and propose to approximate $h(x,t)$ using neural networks, which requires abundance of data and an offline training procedure. In any case, these techniques are computationally costly and subject to approximation errors that may be difficult to quantify.
1.2 Contributions
In this paper we investigate the problem of filtering for Markov state-space models (SSMs) [Sarkka13] under constraints. In a Markov SSM, the state dynamics is described by a sequence of Markov transition kernels $K_{1},K_{2},\ldots,K_{n},\ldots$ and we impose a constraint which is characterized by a sequence of compact subsets ${\mathcal{C}}_{0},{\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{n},\ldots$ of the state space $\mathcal{X}$}. These sets specify the sources and targets of the kernels $K_{n}$, hence they yield a (possibly different) truncation of the state space $\mathcal{X}$} for each $n$. In this setup, we investigate the stability, approximation errors and numerical implementation of the resulting constrained Bayesian filters.
We first prove that, under mild assumptions on the original (unconstrained) SSM, the constrained filter is exponentially stable. Stability is important because it enables the design of numerical filtering algorithms that do not accumulate errors over time [DelMoral01c]. To analyze the approximation error rates of constrained filters, we impose some regularity assumptions both on the SSM and the constraint:
  • (a)
    We parametrize the constraint by an index $l=1,2,\ldots$, in such a way that ${\mathcal{C}}_{n}^{l}\to{\mathcal{X}}$, for every $n$, when $l\to\infty$. Intuitively, the constraints on the state space are relaxed as $l$ increase, and we recover the original (unconstrained) model when $l\to\infty$.
  • (b)
    We assume the compact subsets ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},\ldots$ are sufficiently connected, in the sense that the probability mass allocated to the set ${\mathcal{C}}_{n}^{l}$ by the kernel $K_{n}$ when the initial state is a point ${\mathcal{C}}_{n-1}^{l}$ is bounded away from zero by some constant $\varepsilon_{n}^{l}>0$ that depends on $l$.
Under these assumptions, we find approximation error rates for the constrained filters with respect to $\varepsilon_{n}^{l}$. Moreover, when the positive bounds in assumption (b) above hold uniformly over time, i.e., $\varepsilon_{n}^{l}\geq\varepsilon^{l}>0$ for all $n$, the approximation rates can also be made uniform over $n$. Remarkably, the latter analysis suggests practical guidelines for the selection of the compact sets ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},$... as superlevel sets of the likelihood functions in the unconstrained SSM.
A natural application of the results above is to study the approximation of Bayesian filters by constrained PFs, i.e., PFs that run on the constrained SSM. Under standard regularity assumptions, we obtain time-uniform error rates for PFs that depend both on the number of particles in the Monte Carlo filter and the lower bounds $\varepsilon^{l}>0$ for the transition probabilities from $x\in{\mathcal{C}}_{n-1}^{l}$ to ${\mathcal{C}}_{n}^{l}$ via $K_{n}$. Moreover, we show how the problem of weight degeneracy [Bengtsson08, Snyder08] can be mitigated in the constrained PF when the compacts ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},$... are superlevel sets of the likelihood functions in the unconstrained SSM.
In the continuous-discrete filtering setup, the observations are collected at times $t_{1},t_{2},\ldots,t_{n},$..., and the Markov kernels $K_{n}$ are determined by the SDE that describes the continuous-time state process $\mathtt{X}(t)$ for $t\in[t_{n-1},t_{n}]$. We propose a simple (albeit possibly coarse) approximation of the constrained Markov kernels using so-called barrier functions [Margellos11, Clark21]. These functions can be computed online and they are used to modify the drift of the SDE in order to impose the constraint on the state space. As a result, we obtain a class of guided SDEs that can be easily simulated. Based on this approach, we show some numerical results that illustrate the stability and accuracy of the proposed constrained filters on a stochastic Lorenz 96 model [Grudzien20]
1.3 Organization of the paper
In Section 2 we introduce notation and provide some background on filter stability and constrained SSMs. Section 3 is devoted to the analysis of constrained Bayesian filters, including stability, approximation rates and convergence of constrained PFs. We note that Section 3 is essentially theoretical. Readers interested only in methodological aspects may skip directly to Section 4, where we describe the continuous-discrete filtering setup, discuss Doob’s $h$-transform for guided SDEs, and then introduce the proposed barrier-constrained SDEs. In Section 5 we present numerical results for selected constrained Bayesian filters (using the barrier approach) and the stochastic Lorenz 96 model. Finally, Section 6 is devoted to the conclusions.
2 Background
2.1 State space models and Bayesian filtering
Let us construct a sequence of $d_{x}$-dimensional random vectors $\{X_{n}\}_{n\geq 0}$ on the probability space $(\Omega,{\mathcal{F}},\mathbb{P})$. The law of $X_{0}$ is denoted $\pi_{0}$ and the (random) dynamics of $X_{n}$ are determined by a sequence of Markov kernels
(1) $K_{n}:{\mathcal{X}}\times{\mathcal{B}}({\mathcal{X}})\mapsto[0,1],\quad n>0,$
where ${\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}$ and ${\mathcal{B}}({\mathcal{X}})$ is the Borel $\sigma$-algebra of subsets of $\mathcal{X}$}. We also construct a sequence of uniformly bounded potential functions
(2) $g_{n}:{\mathcal{X}}\mapsto(0,\|g\|_{\infty}],$
where $\|g\|_{\infty}=\sup_{x\in{\mathcal{X}},n\in\mathbb{N}}|g_{n}(x)|<\infty$.
Definition 2.1.
If $K=\{K_{n}\}_{n\geq 1}$ is the sequence of Markov kernels and $g=\{g_{n}\}_{n\geq 1}$ is the sequence of bounded potentials, then we can succinctly denote the system of interest as ${\mathcal{S}}:=(\pi_{0},K,g)$ and we refer to $\mathcal{S}$} as a Markov state space model (SSM).
Let ${\mathcal{P}}({\mathcal{X}})$ denote the family of probability measures on $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ and, for any real function $f:{\mathcal{X}}\mapsto\mathbb{R}$ and $\mu\in{\mathcal{P}}({\mathcal{X}})$, let $\mu(f)$ be shorthand for $\int_{\mathcal{X}}f{\sf d}\mu$. We can define the sequence of operators $\Phi_{n}:{\mathcal{P}}({\mathcal{X}})\mapsto{\mathcal{P}}({\mathcal{X}}),$
(3) $\Phi_{n}(\mu)(f):=\frac{K_{n}\mu(fg_{n})}{K_{n}\mu(g_{n})}$
where $K_{n}\mu({\sf d}z):=\int_{{\mathcal{X}}}K_{n}(x,{\sf d}z)\mu({\sf d}x)$. If we let $\pi_{0}$ be the probability law of the initial state $X_{0}$, then the operators $\{\Phi_{n}\}_{n\geq 1}$ yield the sequence of probability measures
(4) $\pi_{n}:=\Phi_{n}(\pi_{n-1}),\quad n\geq 1,$
generated by the model $\mathcal{S}$}. The composition of consecutive operators is denoted
$\Phi_{m:n}=\Phi_{n}\circ\Phi_{n-1}\circ\cdots\circ\Phi_{m+1},$
for $m\leq n$, in such a way that $\pi_{n}=\Phi_{m:n}(\pi_{m})$. We adopt the convention $\Phi_{n:n}(\pi)=\pi$. Also note that $\Phi_{n-1:n}=\Phi_{n}$.
In Bayesian filtering applications, the state sequence $\{X_{n}\}_{n\geq 0}$ is related to an observation sequence $\{Y_{n}\}_{n\geq 1}$, where $Y_{n}\in{\mathcal{Y}}\subseteq\mathbb{R}^{d_{y}}$, and $\mathcal{Y}$} is a $d_{y}$-dimensional observation space. In this setting, the potential $g_{n}(x)$ is the likelihood of the state $X_{n}=x_{n}$ given the observation $Y_{n}=y_{n}$, and $\pi_{n}$ is the conditional probability law of $X_{n}$ given $Y_{1}=y_{1},\ldots,Y_{n}=y_{n}$, often referred to as the optimal (Bayesian) filter at time $n$. Also, the measure
(5) $\xi_{n}({\sf d}x):=K_{n}\pi_{n-1}({\sf d}x)=\int K_{n}(x^{\prime},{\sf d}x)\pi_{n-1}({\sf d}x^{\prime})$
is the conditional probability law of $X_{n}$ given $Y_{1}=y_{1},\ldots,Y_{n-1}=y_{n-1}$, often referred to as the (one step ahead) predictive distribution of $X_{n}$. Bayes’ theorem readily yields the relationship
(6) $\pi_{n}({\sf d}x)=\frac{g_{n}(x)\xi_{n}({\sf d}x)}{\xi_{n}(g_{n})}.$
Hereafter, we refer to $\Phi_{n}$ as a prediction-update (PU) operator, as it comprises the one-step-ahead prediction of (5) and the Bayesian update of (6). We also employ the term likelihood for $g_{n}$, rather than potential, in order to emphasize the relationship with the observations in the filtering problem.
2.2 A topology on the set of state space models
Let $\mathbb{S}$ be the set of SSMs of the form ${\mathcal{S}}=(\pi_{0},K,g)$ for which the sequence $\{\pi_{n}\}_{n\geq 1}$ is well defined. In particular, we assume that $K_{n}\pi_{n-1}(g_{n})>0$ for all $n\geq 1$, in order to guarantee that $\Phi_{n}(\pi_{n-1})=\pi_{n}\in{\mathcal{P}}({\mathcal{X}}_{n})$ is well defined.
For any two measures $\alpha,\beta\in{\mathcal{P}}({\mathcal{X}})$, let
$\left\|\alpha-\beta\right\|_{\rm tv}:=\sup_{A}|\alpha(A)-\beta(A)|,$
where the supremum is taken over all measurable subsets of $\mathcal{X}$}, denote the total variation norm of the difference $\alpha-\beta$. Following [Crisan20], we impose a topology $\mathbb{D}$ on the set of SSMs $\mathbb{S}$. To be specific, take some ${\mathcal{S}}\in\mathbb{S}$ and let ${\mathcal{S}}^{l}=(\pi_{0}^{l},K^{l},g^{l})$, $l\geq 1$, be a sequence is $\mathbb{S}$, with $K^{l}=\{K_{n}^{l}\}_{n\geq 1}$ and $g^{l}=\{g_{n}^{l}\}_{l\geq 1}$. We say that ${\mathcal{S}}^{l}$ converges to $\mathcal{S}$}, and denote
$\lim_{l\to\infty}{\mathcal{S}}^{l}={\mathcal{S}},$
when
$\lim_{l\to\infty}\left\|\pi_{0}-\pi_{0}^{l}\right\|_{\rm tv}$ $=$ $0,$
$\lim_{l\to\infty}\left\|K_{n}(x,\cdot)-K_{n}(x,\cdot)\right\|_{\rm tv}$ $=$ $0~~\text{for every $n\geq 1$, and}$
$\lim_{l\to\infty}g_{n}^{l}(x)$ $=$ $g_{n}(x)~~\text{for every $n\geq 1$ and $x\in{\mathcal{X}}$.}$
Let $\{\Phi_{n}^{l}\}_{n\geq 1}$ denote the PU operators generated by model ${\mathcal{S}}^{l}$, which in turn yield the sequence of measures $\pi_{n}^{l}=\Phi_{0:n}(\pi_{0}^{l})$. The topology $\mathbb{D}$ has the property that convergence of the sequence of models ${\mathcal{S}}^{l}\to{\mathcal{S}}$ implies convergence of the probability laws $\pi_{n}^{l}\to\pi_{n}$ in total variation. We have the following lemma taken from [Crisan20].
Lemma 2.2.
Let ${\mathcal{S}}^{l}=(\pi_{0}^{l},K^{l},g^{l})$, $l\geq 0$, and ${\mathcal{S}}=(\pi_{0},K,g)$ be elements of $\mathbb{S}$ with corresponding PU operators $\Phi_{n}^{l}$ and $\Phi_{n}$, $n\geq 1$, respectively. If $\lim_{l\to\infty}{\mathcal{S}}^{l}={\mathcal{S}}$, then $\lim_{l\to\infty}\left\|\Phi_{0:n}^{l}(\pi_{0}^{l})-\Phi_{0:n}(\pi_{0})\right\|_{\rm tv}=0$.
2.3 Markov kernels and stability of the optimal filter
Let $\{\Phi_{n}\}_{n\geq 0}$ be the sequence of PU operators generated by the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the likelihoods $g=\{g_{n}\}_{n\geq 1}$ according to (3).
Definition 2.3.
The sequence of optimal filters $\{\pi_{n}\}_{n\geq 1}$, where $\pi_{n}=\Phi_{n}(\pi_{n-1})$, is stable when $\lim_{n\to\infty}\left\|\Phi_{0:n}(\mu)-\Phi_{0:n}(\eta)\right\|_{\rm tv}=0$ for any pair of probability measures $\eta,\mu\in{\mathcal{P}}({\mathcal{X}})$.
The terminology ‘stability of the optimal filter’ is well established in the literature [DelMoral01c, DelMoral04]. We note, however, that stability is a property of the operators $\Phi_{n}$, rather than the probability laws $\pi_{n}$ (see also [Crisan20] for further details).
A sufficient condition for $\Phi_{n}$ to be stable is that the Markov kernels in $K$ are mixing and the likelihood functions in $g$ integrable w.r.t. some reference probability density functions (pdfs) $u=\{u_{n}\}_{n\geq 1}$ on $\mathcal{X}$}.
Definition 2.4.
The Markov kernels $K=\{K_{n}\}_{n\geq 1}$ are mixing with constant $0<\gamma<1$ when
  • (a)
    $K_{n}(x^{\prime},{\sf d}x)=k_{n}(x^{\prime},x){\sf d}x$ for every $n\geq 1$, where $k_{n}:{\mathcal{X}}\times{\mathcal{X}}\mapsto[0,\infty)$ is a pdf w.r.t. the Lebesgue measure ${\sf d}x$, and
  • (b)
    there is a sequence of pdfs $u_{n}:{\mathcal{X}}\mapsto[0,\infty)$ such that
    $\gamma u_{n}(x)\leq k_{n}(x^{\prime},x)\leq\frac{1}{\gamma}u_{n}(x)\quad\text{for every $n\geq 1$ and every $x^{\prime}\in{\mathcal{X}}$.}$
Lemma 2.5.
Let $\{\Phi_{n}\}_{n\geq 1}$ be the PU operators generated by the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the potential functions $g=\{g_{n}\}_{n\geq 1}$. Assume that the Markov kernels are mixing with constant $\gamma>0$ and pdfs $u=\{u_{n}\}_{n\geq 1}$, and
$\int_{\mathcal{X}}g_{n}(x)u_{n}(x){\sf d}x<\infty,\quad\text{for all $n\geq 1$}.$
Then, for any $n\geq m\geq 0$, the composition of PU operators satisfies the inequality
$\left\|\Phi_{m:n}(\mu)-\Phi_{m:n}(\eta)\right\|_{\rm tv}\leq\frac{(1-\gamma^{2})^{n-m}}{\gamma^{2}}\left\|\mu-\eta\right\|_{\rm tv}$
and, in particular, $\lim_{n\to\infty}\left\|\Phi_{0:n}(\mu)-\Phi_{0:n}(\eta)\right\|_{\rm tv}=0$ for any $\mu,\eta\in{\mathcal{P}}({\mathcal{X}})$.
This is a well known result (see Lemma 10 in [Kunsch05]). The mixing condition in Definition 2.4 is sufficient but not necessary for stability.
2.4 Particle filters
A particle filter is a recursive Monte Carlo method to approximate the sequence of measures $\{\pi_{n}\}_{n\geq 0}$. Its simplest version, often termed the bootstrap filter or the standard particle filter [Gordon93, Kitagawa96] (see also [Liu98, Doucet00, Djuric03]), is outlined in Algorithm 1. The Monte Carlo estimate of $\pi_{n}$ is denoted $\pi_{n}^{N}$ and can be written as
(7) $\pi_{n}^{N}({\sf d}x)=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{n}^{i}}({\sf d}x),$
where $\delta_{x}$ is the Dirac delta measure located at $x\in{\mathcal{X}}_{n}$ and $x_{n}^{1},\ldots,x_{n}^{N}$ are Monte Carlo samples from the state space $\mathcal{X}$}.
Algorithm 1 The standard particle filter for a SSM ${\mathcal{S}}=(\pi_{0},K,g)$.
  • Inputs: model ${\mathcal{S}}=(\pi_{0},K,g)$ and number of particles $N$.
    Outputs: approximation $\pi_{n}^{N}$ of the optimal filter $\pi_{n}$ at every step $n\geq 1$.
  • Initialisation. At time $n=0$, draw $N$ iid samples $x_{0}^{i}\sim\pi_{0}$, $i=1,\ldots,N$.
  • Recursive step. At time $n\geq 1$:
    1. 1.
      For $i=1,\ldots,N$,
      • draw $\bar{x}_{n}^{i}\sim K_{n}(x_{n-1},{\sf d}x)$ and
      • compute $\bar{w}_{n}^{i}=g_{n}\left(\bar{x}_{n}^{i}\right)$.
    2. 2.
      Normalise the weights, $w_{n}^{i}=\frac{\bar{w}_{n}^{i}}{\sum_{j=1}^{N}\bar{w}_{n}^{j}}$, $i=1,\ldots,N$.
    3. 3.
      Resample: let $\bar{\pi}_{n}^{N}({\sf d}x)=\sum_{j=1}^{N}w_{n}^{j}\delta_{\bar{x}_{n}^{j}}({\sf d}x)$ and draw $N$ iid samples $x_{n}^{i}\sim\bar{\pi}_{n}^{N}$, $i=1,\ldots,N$.
If the Markov SSM of interest, ${\mathcal{S}}=(\pi_{0},K,g)$, yields a stable optimal filter, then it can be proved that the particle filter approximation $\pi_{n}^{N}$ converges uniformly over time towards the optimal filter $\pi_{n}$. In particular, one has the following result.
Theorem 2.6.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a Markov SSM and assume that
  • (a)
    the kernels $K=\{K_{n}\}_{n\geq 1}$ are mixing with constant $\gamma>0$, and
  • (b)
    there are uniform bounds $\inf_{n\geq 1,x\in{\mathcal{X}}}g_{n}(x)>\zeta>0$ and $\|g\|_{\infty}:=\sup_{n\geq 1}\|g_{n}\|_{\infty}<\infty$ on the potentials $g=\{g_{n}\}_{n\geq 1}$
Then, for any measurable $f:{\mathcal{X}}\mapsto\mathbb{R}$ with $\|f\|\leq 1$, there is a constant $c_{p}<\infty$, independent of $n$, NN, $f$, $d_{x}$ and $d_{y}$, such that
$\sup_{n\geq 1}\|\pi_{n}(f)-\pi_{n}^{N}(f)\|_{p}\leq\frac{c_{p}\|g\|_{\infty}}{\zeta\gamma^{6}}\times\frac{1}{\sqrt{N}}\quad\text{for all $p\geq 1$.}.$
Uniform convergence of the PF is also well known and similar results can be found in [DelMoral04, Kunsch05, Whiteley13b]. We simply note that the mixing assumption (a) in Theorem 2.6 is the same as in [Kunsch05, Theorem 2] (which states uniform convergence in probability of $\pi_{n}^{N}\to\pi_{n}$), and slightly weaker than the usual mixing assumption in [DelMoral04, Theorem 7.4.4]. An explicit proof is presented in Appendix A.
3 Filtering on constrained state space models
3.1 Constrained models
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a Markov SSM in $\mathbb{S}$ and choose a sequence ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ of compact subsets of the state space $\mathcal{X}$}. We consider a constrained SSM $\hat{\mathcal{S}}=(\hat{\pi}_{0},\hat{K},g)\in\mathbb{S}$, where $\hat{K}=\{\hat{K}_{n}\}_{n\geq 1}$,
(8) $\hat{K}_{n}(x^{\prime},{\sf d}x):=\frac{\mathbbm{1}_{{\mathcal{C}}_{n}}(x)K_{n}(x^{\prime},{\sf d}x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x^{\prime},{\sf d}x)},$
$\mathbbm{1}_{{\mathcal{C}}_{n}}$ is the indicator function for the set in the subscript, and $\hat{\pi}_{0}({\sf d}x)=\frac{\mathbbm{1}_{{\mathcal{C}}_{0}}(x)\pi_{0}({\sf d}x)}{\int_{{\mathcal{C}}_{0}}\pi_{0}({\sf d}x)}$. It is easy to see that $\hat{K}_{n}(x^{\prime},{\mathcal{X}})=1$ for any $x^{\prime}\in{\mathcal{X}}$ and hence the kernels in $\hat{K}$ are well defined whenever the original kernels in $K$ are well defined.
The PU operator for model $\hat{\mathcal{S}}$ is constructed as in Eq. (3), i.e., for a real, measurable test function $f:{\mathcal{X}}\mapsto\mathbb{R}$ we have
(9) $\hat{\Phi}_{n}(\mu)(f):=\frac{\hat{K}_{n}\mu(fg_{n})}{\hat{K}_{n}\mu(g_{n})}$
and, given an initial distribution $\hat{\pi}_{0}$, it yields the sequence of marginal posterior distributions $\hat{\pi}_{n}=\hat{\Phi}_{n}(\hat{\pi}_{n-1})=\hat{\Phi}_{0:n}(\hat{\pi}_{0})$, where $\hat{\Phi}_{0:n}=\hat{\Phi}_{n}\circ\cdots\circ\hat{\Phi}_{1}$. Also note that
$\hat{\pi}_{n}({\sf d}x)=\hat{\Phi}_{n}(\hat{\pi}_{n-1})({\sf d}x)=\frac{g_{n}(x)\hat{\xi}_{n}({\sf d}x)}{\hat{\xi}_{n}(g_{n})},$
where $\hat{\xi}_{n}({\sf d}x)=\hat{K}_{n}\hat{\pi}_{n-1}({\sf d}x)$ is the constrained one-step-ahead predictive measure.
3.2 Stability
Let us assume that the SSM of interest, $\mathcal{S}$}, satisfies some mild regularity conditions regarding the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the potentials $g=\{g_{n}\}_{n\geq 1}$.
{assumption}
The elements of model ${\mathcal{S}}=(\pi_{0},K,g)\in\mathbb{S}$ satisfy the conditions below:
  • there are density functions $k_{n}:{\mathcal{X}}\times{\mathcal{X}}\mapsto(0,\infty)$ such that $K_{n}(x^{\prime},{\sf d}x)=k_{n}(x^{\prime},x){\sf d}x$ (in particular, $k_{n}(x^{\prime},x)>0$ for any $x^{\prime},x\in{\mathcal{X}}$ and $n\geq 1$),
  • $\|k\|_{\infty}:=\sup_{n,x^{\prime},x}k_{n}(x^{\prime},x)<\infty$, and
  • $g_{n}>0$ for every $n\geq 1$ and $\|g\|_{\infty}:=\sup_{n\geq 1}\|g_{n}\|_{\infty}<\infty$.
The sufficient conditions for stability in Lemma 2.5 hold most naturally for a constrained SSM $\hat{\mathcal{S}}=(\hat{\pi}_{0},\hat{K},g)$ constructed from a model ${\mathcal{S}}=(\pi_{0},K,g)$ that satisfies Assumption 3.2.
Theorem 3.1.
Let ${\mathcal{S}}=(\pi_{0},K,g)\in\mathbb{S}$ be a SSM for which Assumption 3.2 holds. If we choose a sequence ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ such that
(10) $\inf_{n\geq 1}\inf_{(x^{\prime},x)\in{\mathcal{C}}_{n-1}\times{\mathcal{C}}_{n}}k_{n}(x^{\prime},x)\geq\bar{\gamma}$
for some $\bar{\gamma}>0$, then there is some constant $\gamma>0$ such that
(11) $\left\|\hat{\Phi}_{0:n}(\mu)-\hat{\Phi}_{0:n}(\eta)\right\|_{\rm tv}\leq\frac{(1-\gamma^{2})^{n}}{\gamma^{2}}\left\|\mu-\eta\right\|_{\rm tv},$
where $\hat{\Phi}_{n}$ is the constrained PU operator in (9). In particular,
$\lim_{n\to\infty}\left\|\hat{\Phi}_{0:n}(\mu)-\hat{\Phi}_{0:n}(\eta)\right\|_{\rm tv}=0\quad\text{for any $\mu,\eta\in{\mathcal{P}}({\mathcal{X}})$}$
and the SSM $\hat{\mathcal{S}}$ generates a stable sequence of marginal probability laws $\{\hat{\pi}_{n}\}_{n\geq 0}$.
Proof 3.2.
If (10) holds then, for any $x_{o}\in{\mathcal{C}}_{n-1}$, we have
(12) $0<\frac{\bar{\gamma}}{\|k\|_{\infty}}<\frac{k_{n}(x^{\prime},x)}{k_{n}(x_{o},x)}<\frac{\|k\|_{\infty}}{\bar{\gamma}}<\infty,\quad\forall(x^{\prime},x)\in{\mathcal{C}}_{n-1}\times{\mathcal{C}}_{n}$
where the densities $k_{n}(x^{\prime},x)$ and $k_{n}(x_{o},x)$ are not normalised on ${\mathcal{C}}_{n}$. Moreover, (10) also yields
$\bar{\gamma}|{\mathcal{C}}_{n}|\leq\int_{{\mathcal{C}}_{n}}k_{n}(x_{o},x){\sf d}x\leq\|k\|_{\infty}|{\mathcal{C}}_{n}|,\quad\text{for any $x_{o}\in{\mathcal{C}}_{n-1}$,}$
where $|{\mathcal{C}}_{n}|=\int_{{\mathcal{C}}_{n}}{\sf d}x$. As a consequence,
(13) $\frac{\bar{\gamma}}{\|k\|_{\infty}}\leq\frac{\int_{{\mathcal{C}}_{n}}k_{n}(x^{\prime},x){\sf d}x}{\int_{{\mathcal{C}}_{n}}k_{n}(x_{o},x){\sf d}x}\leq\frac{\|k\|_{\infty}}{\bar{\gamma}}$
and combining (12) with (13) yields
(14) $\frac{\bar{\gamma}^{2}}{\|k\|_{\infty}^{2}}<\frac{\hat{k}_{n}(x^{\prime},x)}{u_{n}(x)}<\frac{\|k\|_{\infty}^{2}}{\bar{\gamma}^{2}}<\infty,\quad\forall(x^{\prime},x)\in{\mathcal{C}}_{n-1}\times{\mathcal{C}}_{n},$
where $\hat{k}_{n}(x^{\prime},x)=\frac{k_{n}(x^{\prime},x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x^{\prime},{\sf d}x)}$ and $u_{n}(x)=\frac{k_{n}(x_{o},x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x_{o},{\sf d}x)}$ are normalised probability densities on ${\mathcal{C}}_{n}$. The inequalities in (14) imply that the constrained kernels $\hat{K}=\{\hat{K}_{n}\}_{n\geq 1}$ are mixing with constant $\gamma:=\frac{\bar{\gamma}^{2}}{\|k\|_{\infty}^{2}}>0$ and the pdfs $u_{n}(x)$ satisfy $\int_{{\mathcal{C}}_{n}}g_{n}(x)u_{n}(x){\sf d}x<\|g\|_{\infty}<\infty$.
Therefore, we can readily apply Lemma 2.5 to obtain the desired inequality (11).
If the state $X_{n}$ is a homogeneous process, i.e., $K_{n}=K$ for some Markov kernel $K:{\mathcal{X}}\times{\mathcal{B}}({\mathcal{X}})\mapsto[0,1]$, and Assumption 3.2 holds, it is always possible to choose compact sets $\{{\mathcal{C}}_{n}\}_{n\geq 1}$ such that the uniform lower bound in (10) exists. Also note that, while Gaussian kernels are not mixing according to Definition 2.4, they satisfy the conditions in Assumption 3.2 and, therefore, they can be constrained to yield a stable sequence of posterior measures $\hat{\pi}_{n}$, $n\geq 0$.
3.3 Approximation errors
We are interested in constructing constrained models $\hat{\mathcal{S}}$ for which the posterior marginals $\hat{\pi}_{n}$ are close to the marginals $\pi_{n}$ generated by the original model $\mathcal{S}$} under (reasonable) regularity assumptions. To this end, we adopt the framework of [Crisan20]. In particular, consider the sequence $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ where each element ${\mathcal{C}}^{l}=\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0}$ is a family of compact sets that can be used to constrain the model $\mathcal{S}$}. Specifically, we construct the sequence of constrained models $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},\hat{K}^{l},g)$, where $\hat{K}^{l}=\{\hat{K}_{n}^{l}\}_{n\geq 1}$,
$\hat{K}_{n}^{l}(x^{\prime},{\sf d}x)=\frac{\mathbbm{1}_{{\mathcal{C}}_{n}^{l}}(x)K_{n}(x^{\prime},{\sf d}x)}{\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)}\quad\text{and}\quad\hat{\pi}_{0}^{l}({\sf d}x)=\frac{\mathbbm{1}_{{\mathcal{C}}_{0}^{l}}(x)\pi_{0}({\sf d}x)}{\int_{{\mathcal{C}}_{0}^{l}}\pi_{0}({\sf d}x)}.$
If ${\mathcal{C}}^{l}$ is constructed in such a way that $\lim_{l\to\infty}{\mathcal{C}}_{n}^{l}={\mathcal{X}}$ for every $n$, then it is apparent that $\lim_{l\to\infty}\hat{\mathcal{S}}^{l}={\mathcal{S}}$ in the topology $\mathbb{D}$ and, by Lemma 2.2,
$\lim_{l\to\infty}\hat{\pi}_{n}^{l}=\lim_{l\to\infty}\hat{\Phi}_{0:n}^{l}(\hat{\pi}_{0}^{l})=\pi_{n}\quad\text{in total variation,}$
where the operators $\hat{\Phi}_{n}^{l}$, $n\geq 1$, are constructed in the obvious way given the kernel $\hat{K}_{n}^{l}$ and Eq. (9). We also denote $\hat{\xi}_{n}^{l}({\sf d}x)=\hat{K}_{n}^{l}\hat{\pi}_{n-1}^{l}({\sf d}x)$.
Not every constraint ${\mathcal{C}}^{l}=\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0}$ on the state space $\mathcal{X}$} leads to a “reasonable” approximation of the SSM $\mathcal{S}$} and the Bayesian filters $\{\pi_{n}\}_{n\geq 1}$. In general, the subsets ${\mathcal{C}}_{n}^{l}$ should be chosen in such a way that they capture a significant mass of the transition probability, i.e., that given a point $x^{\prime}\in{\mathcal{C}}_{n-1}^{l}$, the transition probability
$\mathbb{P}(X_{n}\in{\mathcal{C}}_{n}^{l}|X_{n-1}=x^{\prime})=\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)$
should be large enough. An intuitive interpretation of this condition is that the sequence of subsets $\{{\mathcal{C}}_{n}^{l}\}_{n\geq 1}$ should be “in agreement” with the (unconstrained) dynamics of the state sequence $X_{n}$. This is made formal by Assumption 3.3 below.
{assumption}
The family of compact subsets $\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0,l\geq 1}$ of the state space $\mathcal{X}$} is constructed to guarantee that, for every $n\geq 0$, $\lim_{l\to\infty}{\mathcal{C}}_{n}^{l}={\mathcal{X}}$ and there are positive sequences $\{\varepsilon_{n}^{l}\}_{n\geq 0,l\geq 1}$, increasing with $l$, such that
$\pi_{0}({\mathcal{C}}_{0}^{l})\geq\varepsilon_{0}^{l},\quad\inf_{x^{\prime}\in{\mathcal{C}}_{n-1}^{l}}\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)\geq\varepsilon_{n}^{l}\quad\text{and}\quad\lim_{l\to\infty}\varepsilon_{n}^{l}=1~~\forall n\geq 0.$
Based on the assumption above, and mild conditions on the original model $\mathcal{S}$}, we can obtain a first refinement of Lemma 2.2.
Theorem 3.3.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM where $g_{n}>0$ and $\|g_{n}\|_{\infty}<\infty$ for every $n\geq 1$. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} for which Assumption 3.3 holds and let $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},K^{l},g)$, $l\geq 1$, be the sequence of constrained SSMs. For every measurable and real test function $f\in B({\mathcal{X}})$ and every $n\geq 0$ there are finite constants $c_{0},c_{1},\ldots,c_{n}<\infty$, independent of $l$, such that
$\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\leq\sum_{i=0}^{n}c_{i}\frac{1-\varepsilon_{i}^{l}}{\varepsilon_{i}^{l}}.$
See Appendix B for a proof. The terms $\varepsilon_{i}^{l}$, $i\leq n$, are the increasing sequences in Assumption 3.3, which satisfy $\lim_{l\to\infty}\varepsilon_{i}^{l}=1$, hence $\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\to 0$ as expected from Lemma 2.2.
Assumption 3.3 can be strengthened by imposing that $\varepsilon_{n}^{l}\to 1$ uniformly over all $n\geq 0$.
{assumption}
The family of compact subsets $\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0,l\geq 1}$ of the state space $\mathcal{X}$} is constructed to guarantee that, for every $n\geq 0$, $\lim_{l\to\infty}{\mathcal{C}}_{n}^{l}={\mathcal{X}}$ and there is a single increasing sequence $\varepsilon^{l}>0$ such that
$\pi_{0}({\mathcal{C}}_{0}^{l})\geq\varepsilon^{l},\quad\inf_{x^{\prime}\in{\mathcal{C}}_{n-1}^{l}}\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)\geq\varepsilon^{l}\quad\text{and}\quad\lim_{l\to\infty}\varepsilon^{l}=1~~\forall n\geq 0.$
This stronger assumption on the constraints ${\mathcal{C}}=\{{\mathcal{C}}^{l}\}_{l\geq 1}$ yields a straightforward corollary of Theorem 3.3.
Corollary 3.4.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM where $g_{n}>0$ and $\|g_{n}\|_{\infty}<\infty$ for every $n\geq 1$. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} for which Assumption 3.3 holds and let $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},K^{l},g)$, $l\geq 1$, be the sequence of constrained SSMs. For every measurable and real test function $f\in B({\mathcal{X}})$ and every $n\geq 0$ there is a finite constant ${\rm c}_{n}<\infty$ such that
$\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\leq{\rm c}_{n}\frac{1-\varepsilon^{l}}{\varepsilon^{l}}.$
The proof is trivial: Assumption 3.3 yields $\varepsilon_{n}^{l}=\varepsilon^{l}$ for all $n$, and ${\rm c}_{n}=\sum_{i\leq n}c_{i}$, where the $c_{i}$’s are the constants in Theorem 3.3.
The bounds on the approximation errors $|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|$ can be improved based on Assumption 3.3 and additional conditions on the base model $\mathcal{S}$}. Specifically, we may assume that $\mathcal{S}$} generates a stable sequence of optimal filters $\pi_{n}=\Phi_{0:n}(\pi_{0})$. Note that this is possible even if the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ are not mixing according to Definition 2.4. The following assumption identifies models that generate stable filters with arbitrary (possibly non-geometric) contraction rates.
{assumption}
The model ${\mathcal{S}}=(\pi_{0},K,g)\in\mathbb{S}$ yields a stable sequence of optimal filters. In particular, there is a decreasing sequence $\{r_{i}\}_{i\geq 0}$ such that $r_{0}=1$, $\sum_{i=0}^{\infty}r_{i}<\infty$ and
$\left\|\Phi_{m:n}(\mu)-\Phi_{m:n}(\eta)\right\|_{\rm tv}\leq r_{n-m}\left\|\mu-\eta\right\|_{\rm tv}$
for any pair of probability measures $\mu,\eta\in{\mathcal{P}}({\mathcal{X}})$.
Under Assumption 3.3 it is possible to obtain uniform (over time $n$) error bounds on the approximation error $|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|$. However, similar to Theorem 2.6, the attainment of uniform bounds demands some additional conditions related to the likelihoods $g_{n}$. Specifically, we have worked under two alternative assumptions:
  • the likelihoods are uniformly bounded away from 0, i.e., $\inf_{n,x}g_{n}(x)>\zeta>0$, as in the assumptions of Theorem 2.6 (or in Theorem 7.4.4 in [DelMoral04]), or
  • the likelihoods are continuous, vanishing as $\lim_{\|x\|\to\infty}g_{n}(x)=0$, and the compact sets ${\mathcal{C}}_{n}^{l}$ are chosen to be superlevel sets of $g_{n}$ for every $n\geq 1$, i.e., ${\mathcal{C}}_{n}^{l}=\{x\in{\mathcal{X}}:g_{n}(x)\geq\zeta_{n}^{l}\}$ for some real sequences $\zeta_{n}^{l}>0$, with $\lim_{l\to\infty}\zeta_{n}^{l}=0$ for all $n$.
With the first assumption, we have Theorem 3.5 below, while the second assumption yields Theorem 3.6.
Theorem 3.5.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM for which Assumption 3.3 holds. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} that satisfies Assumption 3.3, and let $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},K^{l},g)$, $l\geq 1$, be the resulting sequence of constrained SSMs. If there are constants $\zeta>0$ and $\|g\|_{\infty}<\infty$ such that $\inf_{n\geq 1,x\in{\mathcal{X}}}g_{n}(x)>\zeta$ and $\sup_{n\geq 1,x\in{\mathcal{X}}}g_{n}(x)\leq\|g\|_{\infty}$, respectively, then there is a finite constant $c<\infty$, independent of $n$, such that
(15) $\sup_{|f|\leq 1}\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\leq c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}.$
In particular, $\lim_{l\to\infty}\sup_{n\geq 0}\left\|\pi_{n}-\hat{\pi}_{n}^{l}\right\|_{\rm tv}=0$.
See Appendix C for a proof. Note that the supremum in (15) is taken over all real and measurable test functions $f:{\mathcal{X}}\mapsto\mathbb{R}$ such that $\|f\|_{\infty}\leq 1$.
Theorem 3.6.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM for which Assumption 3.3 holds. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} that satisfies Assumption 3.3, and let $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},K^{l},g)$, $l\geq 1$, be the resulting sequence of constrained SSMs. If, in addition, the likelihoods $\{g_{n}\}_{n\geq 1}$ are continuous, $\lim_{\|x\|\to\infty}g_{n}(x)=0$ for all $n$ and every ${\mathcal{C}}_{n}^{l}\subseteq{\mathcal{X}}$ is a superlevel set of $g_{n}$, then
(16) $\sup_{|f|\leq 1}\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\leq c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}$
for some constant $c<\infty$ independent of $n$. In particular, $\lim_{l\to\infty}\sup_{n\geq 0}\left\|\pi_{n}-\hat{\pi}_{n}^{l}\right\|_{\rm tv}=0$.
See Appendix D for a proof.
We emphasize that the (identical) inequalities (15) and (15) are obtained under different assumptions on the likelihoods $g_{n}$. In Theorem 3.5, we assume that the likelihood $g_{n}(x)$ is bounded away from zero for all $x\in{\mathcal{X}}$, while in Theorem 3.6, however, the assumption is that $g_{n}(x)\to 0$ when $\|x\|\to\infty$. If the observations have the form $Y_{n}=m(X_{n})+U_{n}$, where $m(\cdot)$ is an observation function and $U_{n}$ is Gaussian noise, then a bounded function $m(\cdot)$ yields $g_{n}>\zeta>0$ (and we can apply Theorem 15) whereas an unbounded $m(\cdot)$ typically yields the assumption in Theorem 16.
We also note that, for many models, Assumption 3.3 (which is used in both Theorems 15 and 16) can be hard to reconcile with the the requirement that the compact sets ${\mathcal{C}}_{n}^{l}$ in the constraint be all of them level supersets of their respective likelihoods $g_{n}$. However, the latter condition is appealing because it yields a practical guideline for the construction of constraints and avoids the need for a uniform positive lower bound on the likelihoods (a condition that does not hold for many models). If we impose the weaker Assumption 3.3 on the family of constraints $\{{\mathcal{C}}^{l}\}_{l\geq 1}$, then the error bounds are not uniform over time any more, however we can still have the following result.
Theorem 3.7.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM for which Assumption 3.3 holds. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} that satisfies Assumption 3.3, and let $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},K^{l},g)$, $l\geq 1$, be the resulting sequence of constrained SSMs. If, in addition, the likelihoods $\{g_{n}\}_{n\geq 1}$ are continuous, $\lim_{\|x\|\to\infty}g_{n}=0$ for all $n$ and every ${\mathcal{C}}_{n}^{l}\subseteq{\mathcal{X}}$ is a superlevel set of $g_{n}$, then
$\sup_{|f|\leq 1}\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\leq 2\sum_{i=0}^{n}r_{n-i}\frac{1-\varepsilon_{i}^{l}}{\varepsilon_{i}^{l}}.$
See Appendix E for a proof.
Remark 3.8.
Note that $r_{n}\leq r_{0}=1$, hence this bound is, in general, tighter than the bound in Theorem 3.3, where the constants $c_{i}$ increase exponentially with index $i$. Moreover, if we restrict attention to a finite subsequence $\pi_{0},\ldots,\pi_{n_{0}}$ for some finite $n_{0}$ then we still have a uniform bound $\left\|\pi_{n}-\hat{\pi}_{n}^{l}\right\|_{\rm tv}\leq s_{0}\frac{1-\epsilon_{0}^{l}}{\epsilon_{0}^{l}}$, where $s_{0}=2\sum_{i=0}^{n_{0}}r_{n-i}<\infty$ and $\epsilon_{0}^{l}=\inf_{0\leq i\leq n_{0}}\varepsilon_{i}^{l}$, where $\epsilon_{0}^{l}\to 1$ as $l\to\infty$.
3.4 Constrained particle filters
In this section we investigate to what extent the standard particle filter approximation $\hat{\pi}_{n}^{l,N}$ of the constrained marginal law $\hat{\pi}_{n}^{l}$ can be used as an estimate of the optimal (unconstrained) Bayesian filter $\pi_{n}$. This can be done in different settings, however we adopt the key assumption $\inf_{n\geq 1}g_{n}>\zeta>0$ in order to provide error bounds comparable to those in Theorem 2.6.
Theorem 3.9.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM for which Assumptions 3.2 and 3.3 hold and, in addition, there is a constant $\zeta>0$ such that $\inf_{n\geq 1}g_{n}>\zeta$. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} that satisfies Assumption 3.3 and the inequality
(17) $\inf_{n\geq 1,(x^{\prime},x)\in{\mathcal{C}}_{n-1}^{l}\times{\mathcal{C}}_{n}^{l}}k_{n}(x^{\prime},x)\geq\bar{\gamma}_{l}>0.$
Then, for any $p\geq 1$,
(18) $\sup_{n\geq 1,|f|\leq 1}\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}<c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}+\frac{c_{p}\|g\|_{\infty}}{\zeta\gamma_{l}^{6}\sqrt{N}},$
where $c_{p}<\infty$ is a constant that depends only on $p$. In particular, for every $l<\infty$,
(19) $\lim_{N\to\infty}\sup_{n\geq 1,|f|\leq 1}\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}<c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}$
and there is some constant $\bar{c}<\infty$, independent of $N$ and $n$, such that
(20) $\sup_{n\geq 0,|f|\leq 1}\lim_{l\to\infty}\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}<\frac{\bar{c}}{\sqrt{N}}.$
A triangle inequality yields $\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}\leq\|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\|_{p}+\|\hat{\pi}_{n}^{l}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}$ and we bound the two terms on the right-hand side separately. In particular, the first term on the right-hand-side of inequality (18) follows from Theorem 3.5, while the second term is obtained via Theorem 3.1 (which ensures the exponential stability of the constrained operators $\hat{\Phi}_{n}^{l}$) together with Theorem 2.6 (for the uniform convergence over time of $\hat{\pi}_{n}^{l,N}$ towards $\hat{\pi}_{n}^{l}$). The limit (19) is straightforward, while he bound in (20) can be obtained by the same argument as in the proof of Theorem 2.6. See Appendix F for a complete proof.
3.5 High-dimensional models
The dependence of the potential functions $g_{n}(x)$ on the dimension $d_{y}$ of the observation space $\mathcal{Y}$} has been found to be a major cause for the “collapse of the particle filter” [Bengtsson08, Snyder08] when $d_{x}\geq d_{y}>>1$, as it causes the importance weights to degenerate, meaning that $\max_{1\leq i\leq n}w_{n}^{i}\to 1$ as $d_{y}\to\infty$. This phenomenon can be alleviated within the framework described in Sections 3.3 and 3.4.
To be specific, choose some functions $\zeta_{n}:\mathbb{N}\mapsto(0,\|g_{n}\|_{\infty}]$ and take superlevel sets ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ such that, for every $n\geq 1$, ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:g_{n}(x)\geq\zeta_{n}(d_{y})\}$. Then consider the (normalised) effective sample size (ESS) estimator [Kong94, Doucet00]
$\text{ESS}_{N}:=\frac{1}{N\sum_{i=1}^{N}\left(w_{n}^{i}\right)^{2}}$
which is usually employed to assess whether the weights are evenly distributed or, on the contrary, one (or a few) particles accumulate all the weight. It can be easily seen that $\frac{1}{N}\leq\text{ESS}_{N}\leq 1$: if the particle weights are uniform, $w_{n}^{i}=\frac{1}{N}$ (which would be the case when sampling directly from the optimal filter), then $\text{ESS}_{N}=1$, while $\text{ESS}_{N}=\frac{1}{N}$ when $\max_{1\leq i\leq N}w_{n}^{i}=1$, i.e., there is a single particle that accumulates all the weight. A low (close to $\frac{1}{N}$) $\text{ESS}_{N}$ indicates that the particle filter has degenerated [Kong94, Doucet00] (or collapsed in the terminology of [Snyder08, Bengtsson08]). It can be seen that, for the constrained particle filter,
$\text{ESS}_{N}\geq\frac{\sup_{x\in{\mathcal{C}}_{n}}g_{n}(x)}{\inf_{x\in{\mathcal{C}}_{n}}g_{n}(x)}\geq\frac{\|g_{n}\|_{\infty}}{\zeta_{n}(d_{y})},$
hence the construction of the sets ${\mathcal{C}}_{n}$ in the constraint ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ as superlevel sets of $g_{n}$ provides a way to control degeneracy of the weights (possibly at the cost of an increased approximation error, depending on the model).
4 Continuous-discrete filtering
4.1 State space model
Let us assume that the signal of interest is modelled as a continuous-time Itô process $\mathtt{X}(t)$ in $(\Omega,{\mathcal{F}},\mathbb{P})$, described by the stochastic differential equations (SDEs)
(21) ${\sf d}\mathtt{X}(t)=a(\mathtt{X}(t),t){\sf d}t+\sigma(\mathtt{X}(t),t){\sf d}\mathtt{W}(t),\quad t\in[0,T],$
where $t$ denotes continuous time, $T$ is some finite time horizon, $a:{\mathcal{X}}\times[0,T]\mapsto{\mathcal{X}}$ is the drift function, $\mathtt{W}(t)$ is a standard $d_{w}$-dimensional Wiener process and $\sigma:{\mathcal{X}}\times[0,T]\mapsto\mathbb{R}^{d_{x}\times d_{w}}$ is the diffusion (matrix) coefficient. For a given time grid
(22) $0=t_{0}<t_{1}<\cdots<t_{n}<\cdots<t_{M}=T,$
the discrete-time state sequence is $X_{n}:=\mathtt{X}(t_{n})$, $n\geq 0$, and the Markov kernels are given by
$K_{n}(x^{\prime},{\sf d}x):=\mathbb{P}\left(\mathtt{X}(t_{n})\in{\sf d}x|\mathtt{X}(t_{n-1})=x^{\prime}\right),\quad\text{for $n\geq 1$}.$
One can generate approximate samples from the kernel $K_{n}(x^{\prime},{\sf d}x)$ using some numerical scheme for the SDE (21) on a possibly refined time grid
$t_{n,0}^{\prime}=t_{n-1}<t_{n,1}^{\prime}<\ldots<t_{n,j}^{\prime}<\cdots<t_{n,J}^{\prime}=t_{n},$
with initial condition $X_{n-1}=\mathtt{X}(t_{n-1})=x^{\prime}$. For example, a plain Euler-Maruyama scheme yields
$\bar{X}_{n,0}$ $=$ $x^{\prime},\quad\text{and}$
$\bar{X}_{n,j}$ $=$ $\bar{X}_{n,j-1}+\Delta_{n,j}a(\bar{X}_{n,j-1},t_{n,j-1}^{\prime})+\sqrt{\Delta_{n,j}}\sigma(\bar{X}_{n,j-1},t_{n,j-1}^{\prime})Z_{n,j},\quad\text{with~~}Z_{n,j}\sim{\mathcal{N}}(0,I),$
for $j=1,\ldots,J$ and $\Delta_{n,j}=t_{n,j}^{\prime}-t_{n,j-1}^{\prime}$. The last iterate, $\bar{X}_{n,J}$, is an approximate sample from the kernel $K_{n}(x^{\prime},{\sf d}x)$, hence $X_{n}\approx\bar{X}_{n,J}$.
The time grid (22) typically corresponds to the time instants where observations are collected and a likelihood function $g_{n}$ is available. We assume that the observations have the form
(23) $Y_{n}=m(X_{n})+U_{n},\quad n=1,\ldots,M,$
where $U_{n}$ is a sequence of zero-mean, i.i.d., $d_{y}$-dimensional real r.v.s, and $m:{\mathcal{X}}\mapsto{\mathcal{Y}}$ is the measurement or observation function that maps the state space ${\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}$ into the observation space ${\mathcal{Y}}\subseteq\mathbb{R}^{d_{y}}$. The form of the likelihood $g_{n}(x)$ depends on the probability law of the noise variables $U_{n}$ and the observed value $Y_{n}=y_{n}$. For example, if the noise is zero-mean Gaussian with covariance matrix $\Sigma_{u}$, denoted $U_{n}\sim{\mathcal{N}}(0,\Sigma_{u})$, then we let
$g_{n}(x):=\exp\left\{-\frac{1}{2}\left\|y_{n}-m(x)\right\|_{\Sigma_{u}^{-1}}^{2}\right\}$
where $\left\|y_{n}-m(x)\right\|_{\Sigma_{u}^{-1}}^{2}=\left(y_{n}-m(x)\right)^{\top}\Sigma_{u}^{-1}\left(y_{n}-m(x)\right)$ and denotes transposition.
4.2 Constrained kernels
Let ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ be a constraint as described in Section 3.1. Sampling the constrained kernel $\hat{K}_{n}(x^{\prime},dx)$ in (8) is equivalent to simulating paths of the Itô process $\mathtt{X}(t)$ that evolves from $\mathtt{X}(t_{n-1})=x^{\prime}$ up to a point in the set ${\mathcal{C}}_{n}$ at time $t_{n}$, i.e., we simulate $\mathtt{X}(t_{n})$ conditional on the initial condition $\mathtt{X}(t_{n-1})=x^{\prime}$ and the terminal condition $\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}$.
One straightforward approach is to simulate trajectories of (21) and then reject the samples $\mathtt{X}(t_{n})\notin{\mathcal{C}}_{n}$. However this can be extremely inefficient as the dimension $d_{x}$ increases. An alternative approach is to “guide” the Itô process to generate samples contained in ${\mathcal{C}}_{n}$ at time $t_{n}$. It is well known that the Doob $h$-transform [Rogers00]
$h(x,t):=\mathbb{P}\left(\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}\mid\mathtt{X}(t)=x\right),$
which satisfies the backward Kolmogorov equation
$\partial_{t}h+a\cdot\nabla h+\frac{1}{2}\mathrm{tr}(\sigma\sigma^{\top}\nabla^{2}h)=0,\qquad h(x,t_{n})=\mathbbm{1}_{{\mathcal{C}}_{n}}(x),$
can be used for this purpose. To be specific, if we let $\sigma^{2}(\cdot,t):=\sigma(\cdot,t)\sigma(\cdot,t)^{\top}$ for conciseness, then the SDE
(24) ${\sf d}\mathtt{X}^{n}(t)=\left[a(\mathtt{X}^{n}(t),t)+\sigma^{2}(\mathtt{X}^{n}(t),t)\nabla_{x}\log h(\mathtt{X}^{n}(t),t)\right]{\sf d}t+\sigma(\mathtt{X}^{n}(t),t){\sf d}\mathtt{W}(t),\quad\mathtt{X}^{n}(t_{n-1})=x^{\prime},$
describes a modified Itô process $\mathtt{X}^{n}(t)$, with $t\in[t_{n-1},t_{n}]$, such that the law of $\mathtt{X}^{n}(t)$ coincides with the law of $\mathtt{X}(t)$ conditional on $\mathtt{X}(t_{n-1})=x^{\prime}$ and $\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}$. Note, however, that $h(X,t)$ is intractable except for a few particular cases and it is also expensive to approximate.
Let
$S_{n}(t):=\left\{x\in{\mathcal{X}}:h(x,t)\geq\exp\{-r(t)\}\right\},\quad t_{n-1}\leq t\leq t_{n},$
be a flow of superlevel sets of the Doob $h$-transform generated by some function $r:[t_{n-1},t_{n}]\mapsto(0,\infty)$, such that $\lim_{t\to t_{n}}r(t)=\infty$ in order to ensure $S_{n}(t_{n})={\mathcal{C}}_{n}$. Given $r(t)$, there is some $\epsilon_{r}\geq 0$ such that
$\mathbb{P}\left(\mathtt{X}^{n}(t)\in S_{n}(t)\right)\geq 1-\epsilon_{r},\quad\forall t\in[t_{n-1},t_{n}].$
Assuming $\epsilon_{r}<<1$, the deterministic flow $S_{n}(t)$ captures most of the probability mass of the constrained process $\mathtt{X}^{n}(t)$ in Eq. (24). Intuitively, we can interpret it as a “high probability tube” connecting the initial condition $\mathtt{X}(t_{n-1})=x^{\prime}$ with the terminal set ${\mathcal{C}}_{n}$ specified by the constraint $\mathcal{C}$}.
4.3 Barrier-constrained kernels
While approximating the Doob $h$-transform $h(x,t)$ is possible (several recent papers [Chopin23, Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25] have tackled this computation, e.g., using local linearisation techniques or neural networks), we explore a simpler alternative where we construct a modified SDE that mimics the structure of (24) but replaces $\log h(x,t)$ by a simple-to-evaluate, differentiable “barrier function” $\log b(x,t)$. This barrier is designed to
  • penalize the continuous-time state trajectories that depart from a high probability tube denoted $\hat{S}_{n}(t)\subset{\mathcal{X}}$, and
  • leave trajectories $\mathtt{X}^{n}(t)\in\hat{S}_{n}(t)$ unaltered.
The tube $\hat{S}_{n}(t)$ is the deterministic flow of subsets of $\mathcal{X}$} constructed by interpolating the constraint sets ${\mathcal{C}}_{n-1}$ and ${\mathcal{C}}_{n}$. In particular, $\hat{S}_{n}(t)$ is defined for $t\in[t_{n-1},t_{n}]$, with initial value $\hat{S}_{n}(t_{n-1})={\mathcal{C}}_{n-1}$ and terminal value $\hat{S}_{n}(t_{n})={\mathcal{C}}_{n}$. Again, this construction mimics the flow $S_{n}(t)$ of high-probability sets described in Section 4.2 above. Specifically, the proposed barrier-constrained SDE takes the form
(25) ${\sf d}\mathtt{X}^{n}(t)=\left[a(\mathtt{X}^{n}(t),t)-\mu\sigma^{2}(\mathtt{X}^{n}(t),t)\nabla_{x}\log b(\mathtt{X}^{n}(t),t)\right]{\sf d}t+\sigma(\mathtt{X}^{n}(t),t){\sf d}\mathtt{W}(t),\quad\mathtt{X}^{n}(t_{n-1})=x^{\prime},$
where $\mu>0$ is a parameter that controls the strength of the penalization when $\mathtt{X}^{n}(t)\notin\hat{S}_{n}(t)$.
Let ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:\log g_{n}(x)\geq\log\upsilon_{n}\}$ be a superlevel set of the likelihood $g_{n}$, in agreement with the assumptions of Theorem 3.6. This construction implies that the sets $\hat{S}_{n}(t)$ depend on the log-likelihood $g_{n}$, hence it is useful to make the barrier depend on this quantity as well. This is relatively simple and can be done in a number of ways. In this paper we use a shifted soft-plus function with parameters $\varrho$ and $\kappa\geq 1$,
$u_{\varrho,\kappa}(z):=\frac{1}{\kappa}\log\left(1+\exp\{\kappa(z-\varrho)\}\right).$
Note that, for large $\kappa$,
$u_{\varrho,\kappa}(z)\approx\left\{\begin{array}[]{ll}0&\text{when}~~z<\varrho\\ z-\varrho&\text{when}~~z>\varrho\\ \end{array}\right.$.
The barrier function is then constructed as
(26) $\log b(x,t):=u_{\varrho,\kappa}(-\log g^{n}(x,t))$
where $g^{n}(x,t)$ is an interpolated likelihood at time $t\in[t_{n-1},t_{n}]$, satisfying $g^{n}(x,t_{n-1})=g_{n-1}(x)$ and $g^{n}(x,t_{n})=g_{n}(x)$.
In Section 5 below, we illustrate the construction of barrier functions for a Gaussian observation model, and study the numerical performance of some Bayesian filters that rely on the barrier-constrained SDE (25) to approximately simulate from the constrained Markov kernel $\hat{K}_{n}(x^{\prime},{\sf d}x)$.
5 Numerical examples
5.1 Stochastic Lorenz 96 model
To illustrate numerically both the analysis in Section 3 and the methodology described in Section 4, we consider the $d_{x}$-dimensional stochastic Lorenz 96 model with additive noise [Perez-Vieites18, Grudzien20] described by the system of SDEs
(27) ${\sf d}\mathtt{X}_{i}(t)=\left[\mathtt{X}_{i-1}(t)\left(\mathtt{X}_{i+1}(t)-\mathtt{X}_{i-2}(t)\right)-\mathtt{X}_{i}(t)+F\right]dt+\sigma_{x}{\sf d}\mathtt{W}_{i}(t),\quad i=0,\ldots,d_{x}-1,$
with periodic boundaries given by
$\mathtt{X}_{-2}(t)=\mathtt{X}_{d_{x}-2}(t),\quad\mathtt{X}_{-1}(t)=\mathtt{X}_{d_{x}-1}(t),\quad\text{and}\quad\mathtt{X}_{0}(t)=\mathtt{X}_{d_{x}}(t),$
where the forcing constant is $F=8$ (which yields turbulent dynamics), $\mathtt{W}_{i}(t)$, $i=0,\ldots,d_{x}-1$, are standard Wiener processes and $\sigma_{x}>0$ is a constant. The complete state is the $d_{x}\times 1$ vector $\mathtt{X}(t)=\left[\mathtt{X}_{0}(t),\ldots,\mathtt{X}_{d_{x}-1}(t)\right]^{\top}$.
We assume a linear and Gaussian observation model, namely
(28) $Y_{n}=H^{\top}X_{n}+U_{n},$
where $U_{n}\sim{\mathcal{N}}(0,\sigma_{y}^{2}I_{d_{y}})$ and $H$ is a $d_{x}\times d_{y}$ matrix and $X_{n}=\mathtt{X}(t_{n})$. For each independent simulation trial, we generate $H$ randomly as
$H=\left[e_{m_{1}},e_{m_{2}},\ldots,e_{m_{d_{y}}}\right]+V$
where $e_{m}$ is a $d_{x}\times 1$ vector of 0s with a single value of 1 in the $m$-th entry, $V$ is a $d_{x}\times d_{y}$ matrix whose entries are independent Gaussian random variates, namely $V_{i,j}\sim{\mathcal{N}}(0,\sigma_{v}^{2})$ and $\sigma_{v}=5\times 10^{-4}$. The indices are drawn randomly from the set $\{1,\ldots,d_{x}\}$, with equal probabilities and no replacement. Intuitively, the $i$-th observation $Y_{i}$, $i\in\{1,\ldots,d_{y}\}$, corresponds to a linear combination of all the state variables, with $X_{m_{i}}$ being dominant and all other variables causing (individually small) interference in the measurement.
We approximate the ground-truth states $X_{n}$ in our experiments below by generating a discrete-time sequence $\mathtt{X}_{l}\approx\mathtt{X}(t_{l}^{\prime})$, where $t_{l}^{\prime}=l\Delta$, by way of the standard Euler-Maruyama scheme with step size $\Delta=10^{-3}$. We assume that observations are collected every $\Delta_{o}=0.1$ continuous time units. Hence, $X_{n}=\mathtt{X}(n\Delta_{o})\approx\mathtt{X}_{\frac{\Delta_{o}}{\Delta}n}=\mathtt{X}_{100n}$. The Euler-Maruyama method is used in the numerical implementation of all the algorithms in this section.
5.2 Barrier function
Assume arbitrary but fixed observations $Y_{n}=y_{n}$, $n\geq 1$. The likelihood function for the observation model (28) is
$g_{n}(x)=\exp\left\{-\frac{1}{2\sigma_{y}^{2}}\|y_{n}-H^{\top}x\|^{2}\right\}$
where $Y_{n}=y_{n}$ is the observation at time $t_{n}=n\Delta_{o}$. The constraint consists of superlevel sets ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:\log g_{n}(x)\geq\log\upsilon\}$. The tube $\hat{S}_{n}(t)$ is constructed as
$S_{n}(t):=\{x\in{\mathcal{X}}:\log g^{n}(x,t)\geq\log\upsilon\},\quad t\in[t_{n-1},t_{n}],$
where $t_{n}=n\Delta_{o}$ and the functions $g^{n}(x,t)$ are obtained by interpolating the observations $y_{n-1}$ and $y_{n}$. Specifically, we let
$y^{n}(t):=\frac{y_{n}-y_{n-1}}{t_{n}-t_{n-1}}t+\frac{y_{n-1}t_{n}-y_{n}t_{n-1}}{t_{n}-t_{n-1}}~~\text{and}~~g^{n}(x,t):=\exp\left\{-\frac{1}{2\sigma_{y}^{2}}\|y^{n}(t)-H^{\top}x\|^{2}\right\}.$
Then, by construction, $g^{n}(x,t_{n-1})=g_{n-1}(x)$, $g^{n}(x,t_{n})=g_{n}(x)$, $S_{n}(t_{n-1})={\mathcal{C}}_{n-1}$ and $\hat{S}_{n}(t_{n})={\mathcal{C}}_{n}$.
Moreover, we note that
$\hat{S}_{n}(t)=\{x\in{\mathcal{X}}:\log g^{n}(x,t)\geq\log\upsilon\}=\left\{x\in{\mathcal{X}}:\|y^{n}(t)-H^{\top}x\|\leq\sqrt{-\frac{\log\upsilon}{2\sigma_{y}^{2}}}\right\},$
hence, $\hat{S}_{n}(t)$ is the preimage (on the state space $\mathcal{X}$}) of the ball $\{y\in{\mathcal{Y}}:\|y^{n}(t)-y\|\leq R\}$, with radius $R=\sqrt{-\frac{\log\upsilon}{2\sigma_{y}^{2}}}$ (on the observation space $\mathcal{Y}$}). The barrier function in (26) becomes
$\log b(x,t)$ $=$ $u_{\varrho,\kappa}\left(\frac{1}{2\sigma_{y}^{2}}\|y^{n}(t)-H^{\top}x\|^{2}\right)$
$=$ $\frac{1}{\kappa}\log\left(1+\exp\left\{\kappa\left(\frac{1}{2\sigma_{y}^{2}}\|y^{n}(t)-H^{\top}x\|^{2}-\varrho\right)\right\}\right),$
where $\varrho=-\log\upsilon$,
(29) $\nabla_{x}\log b(x,t)=\frac{1}{\sigma_{y}^{2}}s\left(\kappa\left(\frac{1}{2\sigma_{y}^{2}}\|y^{n}(t)-H^{\top}x\|^{2}-\varrho\right)\right)H(H^{\top}x-y^{n}(t)),$
where $s(z)=\frac{1}{1+\exp\{-z\}}$ is the logistic function.
5.3 Stability of constrained filters
Let ${\mathcal{S}}^{A}=(\pi_{0}^{A},K,g)$ and ${\mathcal{S}}^{B}=(\pi_{0}^{B},K,g)$ be two SSMs that differ only on their initial distribution. To assess the stability of the PU operators $\Phi_{n}$ (identical for the two models, as $\Phi_{n}$ depends only on $K_{n}$ and $g_{n}$) we ideally need to evaluate the total variation distance $\left\|\pi_{n}^{A}-\pi_{n}^{B}\right\|_{\rm tv}$, where $\pi_{n}^{A}$ and $\pi_{n}^{B}$ are the filtering distributions at time $n$ generated by model ${\mathcal{S}}^{A}$ and model ${\mathcal{S}}^{B}$, respectively. Unfortunately, this distance is intractable, hence we compute a proxy for $\left\|\pi_{n}^{A}-\pi_{n}^{B}\right\|_{\rm tv}$ by applying a simple discretisation procedure.
Specifically, at the time $t_{n}$ of the $n$-th observation, let us construct the hypercube
${\mathcal{H}}_{n}=\{x\in{\mathcal{X}}:\|x-X_{n}\|_{\infty}\leq r\},$
centered at the true state vector $X_{n}$, with semi-length $r>0$. This hypercube is then uniformly partitioned into $L_{\rm sub}:={\left\lfloor\frac{r}{r_{\text{sub}}}\right\rfloor^{d_{x}}}$ smaller sub-hypercubes ${\mathcal{H}}_{n,l}$, $l=1,\ldots,L_{\rm sub}$, each of semi-length $r_{\rm sub}$. We note that ${\mathcal{H}}_{n}=\cup_{l=1}^{L_{\rm sub}}{\mathcal{H}}_{n,l}$ and ${\mathcal{H}}_{n,l}\cap{\mathcal{H}}_{n,j}=\emptyset$ when $l\neq j$. Given two measures $\alpha_{n},\beta_{n}\in{\mathcal{P}}({\mathcal{X}})$, their total variation distance is approximated as
(30) $\left\|\alpha_{n}-\beta_{n}\right\|_{\rm tv}\approx\frac{1}{2}\sum_{i}\big|\alpha_{n}(\mathcal{H}_{n,i})-\beta_{n}(\mathcal{H}_{n,i})\big|.$
From the SSMs ${\mathcal{S}}^{A}$ and ${\mathcal{S}}^{B}$ we construct constrained models $\hat{\mathcal{S}}^{A}$ and $\hat{\mathcal{S}}^{B}$. The constrained kernels $\hat{K}_{n}$ are implemented using the barrier method of Section 4.3 with parameters $\varrho=2$, $\kappa=1$ and $\mu=10$. We simulate the Lorenz 96 model with dimension $d_{x}=10$, with $d_{y}=6$ observations at every time $t_{n}$, and noise scale parameters $\sigma_{x}^{2}=\frac{1}{2}$ and $\sigma_{y}^{2}=\frac{1}{4}$. The discretisation of the TV distance is carried out with $r=6$ and $r_{\rm sub}=3$. The filters $\hat{\pi}_{n}^{A}$ and $\hat{\pi}_{n}^{B}$ are approximated via standard particle filtering on the constrained models using $N$ particles, and we denote the Monte Carlo approximate filters as $\hat{\pi}_{n}^{A,N}$ and $\hat{\pi}_{n}^{B,N}$, respectively.
Figure 1 illustrates the exponential stability predicted by Theorem 3.1. Specifically, in Figure 1(a), we display the (approximate) total variation distances $\left\|\pi_{n}^{A,N_{1}}-\pi_{n}^{B,N_{1}}\right\|_{\rm tv}$ and $\left\|\pi_{n}^{A,N_{2}}-\pi_{n}^{B,N_{2}}\right\|_{\rm tv}$, obtained with $N_{1}=10,240$ and $N_{2}=1,024,000$ particles, averaged over 128 independent simulation runs. The prior distributions for this experiment are almost mutually singular, namely $\pi_{0}^{A}=\mathcal{N}(X_{0}-1,\dfrac{1}{4}I_{dx})$ and $\pi_{0}^{B}=\mathcal{N}(X_{0}+1,4I_{dx})$. The total variation distances are plotted versus the observation index (discrete time $n$). We see that the total variation distance between the constrained posterior measures with different priors decays exponentially fast, until it hits a Monte Carlo error floor, which is lower as we increase the number of particles in the PF.
In Figure 1(b) we show a numerical estimate of the bound in (11), which is obtained by a nonlinear least squares fit of the constant $\gamma$ (namely, $\hat{\gamma}\approx 0.496$). We see that the bound holds until the total variation curve approaches the Monte Carlo error floor.
((a))
((b))
Figure 1: (a) The distances $\left\|\hat{\pi}_{n}^{A,N_{1}}-\hat{\pi}_{n}^{B,N_{1}}\right\|_{\rm tv}$ and $\left\|\hat{\pi}_{n}^{A,N_{2}}-\hat{\pi}_{n}^{B,N_{2}}\right\|_{\rm tv}$ versus observation index $n$, with number of particles $N_{1}=10,240$ and $N_{2}=1,024,000$ and barrier parameter $\varrho=2$. (b) The distance $\left\|\hat{\pi}_{n}^{A,N}-\hat{\pi}_{n}^{B,N}\right\|_{\rm tv}$ with $N=1,024,000$, barrier parameter $\varrho=2$ and the nonlinear fit $\frac{(1-\hat{\gamma}^{2})^{n}}{\hat{\gamma}^{2}}\left\|\hat{\pi}_{0}^{A}-\hat{\pi}_{0}^{B}\right\|_{\rm tv}$ of the contraction bound in (11).
Next, let us take a sequence of three constraints ${\mathcal{C}}^{l}$, $l=1,2,3$, which differ in the choice of the radius parameter $\varrho$ in the barrier function. In particular, we set $\kappa=1$ and $\mu=10$ for all three constraints, while $\{\varrho_{1},\varrho_{2},\varrho_{3}\}=\{0.5,~~2,~~4\}$. We then repeat the same experiment as in Figure 1, to obtain the decay of the total variation distance versus discrete time $n$.
The results are displayed in Figure 2(a). It illustrates how the volume of the constraint sets ${\mathcal{C}}_{n}^{l}$, determined by the barrier radius $\varrho_{l}$ affects the stability of the constrained filter. As $\varrho_{l}$ decreases, constrained sets become smaller, leading to a stronger contraction of the posterior measures and, hence, a faster decay of the total variation distance. The fitted decay rates are $\{\hat{\gamma}_{1},\hat{\gamma}_{2},\hat{\gamma}_{3}\}=\{0.518,0.505,0.487\}$, where larger $\gamma$ yields a faster decay.
5.4 Volume of constraint sets
In these experiments we study the effect the volume of the constraint sets on the accuracy of the barrier-constrained filters. Similar to the previous experiments we take stochastic Lorenz 96 model with state dimension $d_{x}=10$, $d_{y}=6$ observations at every time $t_{n}$, and noise scale parameters $\sigma_{x}^{2}=\frac{1}{2}$ and $\sigma_{y}^{2}=\frac{1}{4}$. The barrier method is implemented with parameters $\kappa=1$ and $\mu=10$, and radii $\{\varrho_{1},\varrho_{2},\varrho_{3},\varrho_{4}\}=\{0.5,1,2,4\}$. We note again that the radius $\varrho_{l}$ determines the volume of the sets ${\mathcal{C}}_{n}^{l}$. With this setup, we run a standard PF with $N=50,000$ particles for reference, to approximate an “optimal” posterior mean estimator, a standard PF with just $N=500$ particles, and several barrier-constrained PFs (for $\varrho_{1},\ldots,\varrho_{4}$) with $N=500$ particles.
Figure 2(b) displays the results in terms of the normalised mean square error (NMSE) between between the posterior mean estimate $\hat{x}_{n}^{N}=\frac{1}{N}\sum_{i=1}^{N}x_{n}^{i}$ output by each filter and the true state $X_{n}$, namely ${\rm NMSE}_{n}=\frac{T\|X_{n}-\hat{x}_{n}^{N}\|^{2}}{\sum_{m=1}^{T}\|X_{m}\|^{2}}$, where $T$ is the number of observation times. The standard PF with $N=500$ can be expected to lose track of the system state after less than 20 discrete time steps (for $n<20$) while the barrier-constrained filters perform close to the optimal posterior mean estimator for the complete simulation, with $50$ discrete time steps (the results are the same for longer simulations). The NMSE curves have been averaged over 128 independent simulation runs.
((a))
((b))
Figure 2: (a) The (discretised) distance $\left\|\pi_{k}^{A,N}-\pi_{k}^{B,N}\right\|_{\rm tv}$ computed using standard PFs with $N=10,240$ particles and barrier parameters $\{\varrho_{1},\varrho_{2},\varrho_{3}\}=\{0.5,2,4\}$. (b) NMSE versus observation index (discrete time $n$) for the standard PF (with $N=500$ and $N=50,00$ particles) and barrier-constrained PFs with different constraints (parametrised by the barrier radii $\{\varrho_{1},\varrho_{2},\varrho_{3},\varrho_{4}\}=\{0.5,1,2,4\}$.
5.5 Higher dimensional models
In this final experiment we assess the accuracy of the barrier-constrained methods as the dimension of the state ($d_{x}$) and the observations ($d_{y}$) increase. Otherwise, the stochastic Lorenz 96 model parameters are $\sigma_{x}=\sigma_{y}=1$ and the continuous-time horizon is $T=10$, with observations collected every $\Delta_{o}=0.1$ continuous time units. For the barrier-based filters, we set $\mu=50$, $\kappa=100$ and $\varrho=4$.
To be specific, we compare the performance of 5 types of filters:
  • a standard PF (labeled SIR),
  • an auxiliary particle filter (APF) [Pitt01],
  • an ensemble Kalman filter (EnKF) [Evensen03] implemented as in [Perez-Vieites18],
  • a barrier-constrained standard PF (labeled barrier-SIR), and
  • a barrier-constrained EnKF (barrier-EnKF).
In Figure 3(a) we plot the NMSE attained by each filter versus the state dimension $d_{x}$ for a fixed number of particles (or ensemble size for the EnKFs) $N=750$. The observation dimension is set as $d_{y}=\lfloor 0.6d_{x}\rfloor$ and the pattern of observations is fixed for each simulation but chosen randomly across different simulations. The results displayed are an average of 10 independent runs.
We observe that the standard PFs break down quickly as the dimension increases and yield NMSEs over 1, which means that the power of the error is of the same order as the power of the signal $X_{n}$ itself. The performance of the standard EnKF degrades quickly when $d_{x}>N$, while both the barrier-constrained PF and the barrier-constrained EnKF yield stable NMSEs as $d_{x}$ increases. The NMSE attained by the barrier-EnKF filter is around half the error of the barrier-SIR filter.
Figure 3(b) shows the results of a similar experiment where the number of particles for the different algorithms is increased linearly with the model dimension, namely $N=d_{x}$. The simulation shows that the standard PFs still fail to track the stochastic Lorenz 96 model, while the barrier-based algorithms remain stable and improve their performance. The barrier-EnKF filter yields an NMSE which is consistently lower than the NMSE of the standard EnKF for $d_{x}\geq 40$, with a very similar computational cost.
((a))
((b))
Figure 3: NMSE versus state-space dimension $d_{x}$ for SIR, EnKF, APF, barrier-SIR, and barrier-EnKF algorithms. (a) All filters run with a fixed sample size $N=750$. (b) Sample size is scaled with dimension as $N=d_{x}$.
6 Conclusions
We have investigated a class of constrained Bayesian filters for state space models where the source and target of the Markov transition kernels are compact subsets of the state space constructed from the observed data. In this setup, we have analyzed the stability and approximation errors of Bayesian filters. To be specific, we have proved that, under mild assumptions on the original (unconstrained) state-space model, the constrained filter is exponentially stable. Stability is an important feature because it enables the design of numerical filters which do not accumulate error over time. We have also studied the approximation errors of constrained filters and obtained error rates that depend on a parameter of the constraint. Our analysis also provides guidelines on the selection of the sequence of compact sets that form the constraint, namely, that they should be selected as superlevel sets of the likelihood functions in the state space model. The results of the approximation error analysis can also be leveraged to prove that particle filters run on the constrained state space model can be used to approximate the optimal (unconstrained) Bayesian filter under some regularity assumptions. Moreover, the selection of the compacts sets in the constraint as superlevel sets of the likelihood functions provides means to mitigate weight degeneracy in the particle filter.
In the second half of the paper we have turned our attention to the practical implementation of the proposed constrained filters in problems where data are collected over a discrete time grid but the signal of interest is a continuous-time stochastic Itô process characterised by a SDE. We have proposed a relatively simple approximation of the resulting constrained Markov kernels using barrier functions to modify the drift of the original SDE. As a result, we obtain a class of modified SDEs that can be easily simulated. Based on this approach, we have shown some numerical results that illustrate the stability and accuracy of the proposed constrained filters when tracking a (possibly high dimensional) stochastic Lorenz 96 system with partial and noisy observations.
Acknowledgments
JM thanks the mathematical research institute MATRIX, in Australia, where part of this research was performed.
Appendix A Proof of Theorem 2.6
Let $f:{\mathcal{X}}\mapsto\mathbb{R}$ be a measurable test function, with $\|f\|_{\infty}<\infty$. The difference $\pi_{n}(f)-\pi_{n}^{N}(f)$ can be expanded as
$\pi_{n}^{N}(f)-\pi_{n}(f)$ $=$ $\sum_{i=0}^{n-1}\Phi_{n-i:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n-i:n}\left(\Phi_{n-i}\left(\pi_{n-i-1}^{N}\right)\right)(f)$
$+\Phi_{0:n}\left(\pi_{0}^{N}\right)(f)-\Phi_{0:n}\left(\pi_{0}\right)(f).$
and a straightforward application of Minkowski’s inequality yields
(31) $\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}$ $\leq$ $\sum_{i=0}^{n-1}\left\|\Phi_{n-i:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n-i:n}\left(\Phi_{n-i}\left(\pi_{n-i-1}^{N}\right)\right)(f)\right\|_{p}$
$+\left\|\Phi_{0:n}\left(\pi_{0}^{N}\right)(f)-\Phi_{0:n}\left(\pi_{0}\right)(f)\right\|_{p}.$
for any $p\geq 1$.
For $i=0$ in (31), we have
(32) $\left\|\Phi_{n:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n:n}\left(\Phi_{n}\left(\pi_{n-1}^{N}\right)\right)(f)\right\|_{p}=\|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\|_{p}$
and it is straightforward to show that
$\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)$ $=$ $\frac{\xi_{n}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}-\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\pm\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}$
$=$ $\frac{\xi_{n}^{N}(g_{n}f)-K_{n}\pi_{n-1}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}+\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\frac{K_{n}\pi_{n-1}^{N}(g_{n})-\xi_{n}^{N}(g_{n})}{\xi_{n}^{N}(g_{n})}.$
Since $\left|\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\right|=|\Phi_{n}(\pi_{n-1}^{N})(f)|\leq\|f\|_{\infty}$, we readily see that
(33) $\left|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\right|\leq\frac{2\|f\|_{\infty}}{\xi_{n}^{N}(g_{n})}\left|\xi_{n}^{N}(g_{n})-K_{n}\pi_{n-1}^{N}(g_{n})\right|.$
From assumption (b) in the statement of Theorem 2.6 we obtain $\xi_{n}^{N}(g_{n})\geq\zeta>0$ and a straightforward application of the Marcinkiewicz-Zygmund inequality [^1] Note that $K_{n}\pi_{n-1}^{N}({\sf d}x_{n})=\frac{1}{N}\sum_{j=1}^{N}K_{n}(x_{n-1}^{j},{\sf d}x_{n})$ and $\xi_{n}^{N}({\sf d}x_{n})=\frac{1}{N}\sum_{j=1}^{N}\delta_{\bar{x}_{n}^{j}}({\sf d}x_{n})$, where $\bar{x}_{n}^{j}\sim K_{n}(x_{n-1}^{j},{\sf d}x_{n})$. If we let $\bar{\mathcal{F}}_{n-1}$ be the $\sigma$-algebra generated by $\{x_{0:n-1}^{1:N},\bar{x}_{1:n-1}^{1:N}\}$, it is clear that $\mathbb{E}[f(\bar{x}_{n}^{j})|\bar{\mathcal{F}}_{n-1}]=\int f(x_{n})K_{n}(x_{n-1}^{j},{\sf d}x_{n})$ for every j=1,,Nj=1,...,N. yields
(34) $\left\|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\right\|_{p}\leq\frac{2c_{p}\|f\|_{\infty}\|g\|_{\infty}}{\zeta\sqrt{N}},$
where $c_{p}$ is a constant that depends only on $p$ (but is independent of $\xi_{n}^{N}$ and, in particular, of $N$) and we have used the bound $|g_{n}|\leq\|g\|_{\infty}$, also from assumption (b).
For $1\leq i\leq n-1$ in (31), let us first note that, from the mixing assumption (a) in Theorem 2.6 and Lemma 2.5, we arrive at
$\left\|\Phi_{n-i:n}(\pi_{n-i}^{N})-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))\right\|_{\rm tv}$ $\leq$
$\frac{(1-\gamma^{2})^{i-1}}{\gamma^{2}}\left\|\Phi_{n-i+1}(\pi_{n-i}^{N})-\Phi_{n-i+1}(\Phi_{n-i}(\pi_{n-i-1}^{N}))\right\|_{\rm tv}$ $=$
(35) $\frac{(1-\gamma^{2})^{i-1}}{\gamma^{2}}\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{n-i+1}(\pi_{n-i}^{N})(f)-\Phi_{n-i+1}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f).\right|$
However, for any $j\geq 2$, one can show that
(36) $\Phi_{j}(\pi_{j-1}^{N})(f)$ $=$ $\frac{\int f(x_{j})g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})\xi_{j-1}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})\xi_{j-1}^{N}({\sf d}x_{j-1})}$
$=$ $\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
where $K_{j}(g_{j}f)(x_{j-1})=\int f(x_{j})g_{j}(x_{j})K_{j}(x_{j-1},{\sf d}x_{j})$, and, similarly,
(37) $\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)$ $=$ $\frac{\int f(x_{j})g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}$
$=$ $\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}.$
Combining (36) and (37) yields
$\Phi_{j}(\pi_{j-1}^{N})(f)-\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)$ $=$
$\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}-\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}\pm\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$ $=$
$\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)-K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
(38) $+\rho_{j}(f)\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)-\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
where
$\rho_{j}({\sf d}x_{j})=\frac{g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}$
is a probability measure. From (38) and the fact that $|\rho_{j}(f)|\leq\|f\|_{\infty}$ we readily find the inequality
(39) $\left|\Phi_{j}(\pi_{j-1}^{N})(f)-\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)\right|\leq\frac{2\|f\|_{\infty}\left|\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)-K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)\right|}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
and, combining (35) and (39) with j=ni+1j=n-i+1 yields
$\sup_{\|f\|_{\infty}\leq 1}\left|\Phi_{n-i:n}(\pi_{n-i}^{N})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f)\right|$ $\leq$
(40) $\frac{2(1-\gamma^{2})^{i-1}}{\gamma^{2}}\frac{\left|\xi_{n-i}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)-K_{n-i}\pi_{n-i-1}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)\right|}{\xi_{n-i}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)}.$
Recalling assumption (b) of Theorem 2.6 ($\inf_{n}g_{n}\geq\zeta>0$) and the mixing assumption (a) ($\gamma u_{j}(x_{j})\leq k_{j}(x_{j-1},x_{j})\leq\frac{1}{\gamma}u_{j}(x_{j})$) we realise that
$|K_{n-i+1}(g_{n-i+1})g_{n-i}|\leq\frac{\|g\|_{\infty}}{\gamma}\int g_{n-i+1}(x_{n-i+1})u_{n-i+1}(x_{n-i+1}){\sf d}x_{n-i+1}$
and
$K_{n-i+1}(g_{n-i+1})g_{n-i}\geq\gamma\zeta\int g_{n-i+1}(x_{n-i+1})u_{n-i+1}(x_{n-i+1}){\sf d}x_{n-i+1}>0.$
Then, combining (40) with the Marcinkiewicz-Zygmund inequality, applied in the same way as in (34), we readily arrive at the $L^{p}$ bound
(41) $\left\|\Phi_{n-i:n}(\pi_{n-i}^{N})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f)\right\|_{p}\leq\frac{2c_{p}\|g\|_{\infty}(1-\gamma^{2})^{i-1}}{\gamma^{4}\zeta}\times\frac{1}{\sqrt{N}}.$
that holds for any measurable test function $f:{\mathcal{X}}\mapsto\mathbb{R}$ with $\|f\|_{\infty}\leq 1$.
The calculation for the last term on the right hand side of (31) is very similar. Lemma 2.5 yields
(42) $\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{0:n}(\pi_{0}^{N})(f)-\Phi_{0:n}(\pi_{0})(f)\right|\leq\frac{(1-\gamma^{2})^{n-1}}{\gamma^{2}}\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right|$
and
$\left|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right|$ $\leq$
$\left|\frac{K_{1}\pi_{0}^{N}(g_{1}f)-K_{1}\pi_{0}(g_{1}f)}{K_{1}\pi_{0}(g_{1})}+\Phi_{1}(\pi_{0}^{N})(f)\frac{K_{1}\pi_{0}(g_{1})-K_{1}\pi_{0}^{N}(g_{1})}{K_{1}\pi_{0}(g_{1})}\right|$ $\leq$
(43) $\frac{2\|f\|_{\infty}}{K_{1}\pi_{0}(g_{1})}\left|K_{1}\pi_{0}^{N}(g_{1})-K_{1}\pi_{0}(g_{1})\right|$
and, using assumption (b) ($g_{1}>\zeta>0$) and the Marcinkiewicz-Zygmund inequality again, we arrive at
(44) $\left\|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right\|_{p}\leq\frac{2\|f\|_{\infty}\|g\|_{\infty}c_{p}}{\zeta\sqrt{N}},$
where the constant $c_{p}<\infty$ depends only on $p\geq 1$. Hence, for any measurable and real test function $f$ with $\|f\|\leq 1$, the combination of (42), (43) and (44) yields
(45) $\left\|\Phi_{0:n}(\pi_{0}^{N})(f)-\Phi_{0:n}(\pi_{0})(f)\right\|_{p}\leq\frac{2c_{p}\|g\|_{\infty}(1-\gamma^{2})^{n-1}}{\gamma^{2}\zeta\sqrt{N}}.$
Finally, taking together (31), (34), (41) and (45) yields
$\sup_{\|f\|_{\infty}\leq 1}\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}\leq\frac{2c_{p}\|g\|_{\infty}}{\gamma^{6}\zeta}\times\frac{1}{\sqrt{N}}.\quad\Box$
Appendix B Proof of Theorem 3.3
We proceed by induction in time $n$. For $n=0$, we have the constrained prior $\hat{\pi}_{0}^{l}({\sf d}x)=\frac{\mathbbm{1}_{{\mathcal{C}}_{0}^{l}}(x)\pi_{0}({\sf d}x)}{\int_{{\mathcal{C}}_{0}^{l}}\pi_{0}({\sf d}x)}$, hence the approximation error can be written as
$\left|\pi_{0}(f)-\hat{\pi}_{0}^{l}(f)\right|$ $=$ $\left|\int_{\mathcal{X}}f(x)\left(1-\frac{\mathbbm{1}_{{\mathcal{C}}_{0}^{l}}(x)}{\int_{{\mathcal{C}}_{0}^{l}}\pi_{0}({\sf d}x)}\right)\pi_{0}({\sf d}x)\right|$
$=$ $\left|\int_{{\mathcal{C}}_{0}^{l}}f(x)\frac{\int_{{\mathcal{C}}_{0}^{l}}\pi_{0}({\sf d}x)-1}{\int_{{\mathcal{C}}_{0}^{l}}\pi_{0}({\sf d}x)}\pi_{0}({\sf d}x)+\int_{\overline{{\mathcal{C}}_{0}^{l}}}f(x)\pi_{0}({\sf d}x)\right|,$
where the second equality follows from splitting the integral over the partition ${\mathcal{X}}={\mathcal{C}}_{0}^{l}\cup\overline{{\mathcal{C}}_{0}^{l}}$ (we denote $\overline{{\mathcal{C}}_{0}^{l}}={\mathcal{X}}\backslash{\mathcal{C}}_{0}^{l}$) and the fact that $\hat{\pi}_{0}^{l}(\mathbbm{1}_{S})=0$ for any $S\subseteq\overline{{\mathcal{C}}_{0}^{l}}$. Assumption 3.3 together with the equation above yield
(46) $\left|\pi_{0}(f)-\hat{\pi}_{0}^{l}(f)\right|\leq\|f\|_{\infty}\left(\frac{1-\varepsilon_{0}^{l}}{\varepsilon_{0}^{l}}+(1-\varepsilon_{0}^{l})\right)\leq c_{0}\frac{1-\varepsilon_{0}^{l}}{\varepsilon_{0}^{l}},$
where $c_{0}=2\|f\|_{\infty}$ (note that $\varepsilon_{0}^{l}\leq 1$). Inequality (46) is the base case of the induction.
Let us assume that, at time n1n-1, with $n\geq 1$,
(47) $\left|\pi_{n-1}(f)-\hat{\pi}_{n-1}^{l}(f)\right|\leq\sum_{i=0}^{n-1}\tilde{c}_{i}\frac{1-\varepsilon_{i}^{l}}{\varepsilon_{i}^{l}}$
for some constants $\tilde{c}_{i}<\infty$ and the sequences $\varepsilon_{i}^{l}$ in Assumption 3.3, with $0\leq i\leq n-1$.
At time $n$, we first look into the predictive measures $\xi_{n}=K_{n}\pi_{n-1}$ and $\xi_{n}^{l}=\hat{K}_{n}^{l}\hat{\pi}_{n-1}^{l}$. A straightforward triangle inequality yields
(48) $\left|\xi_{n}(f)-\hat{\xi}_{n}^{l}(f)\right|\leq\left|K_{n}\pi_{n-1}(f)-K_{n}\hat{\pi}_{n-1}^{l}(f)\right|+\left|K_{n}\hat{\pi}_{n-1}^{l}(f)-\hat{K}_{n}^{l}\hat{\pi}_{n-1}^{l}(f)\right|.$
If we let $K_{n}(f)(x^{\prime}):=\int_{\mathcal{X}}f(x)K_{n}(x^{\prime},{\sf d}x)$ then the first term on the right-hand side of (48) can be bounded as
(49) $\left|K_{n}\pi_{n-1}(f)-K_{n}\hat{\pi}_{n-1}^{l}(f)\right|=\left|\pi_{n-1}(K_{n}(f))-\hat{\pi}_{n-1}^{l}(K_{n}(f))\right|\leq\|f\|_{\infty}\sum_{i=0}^{n-1}\tilde{c}_{i}\frac{1-\varepsilon_{i}^{l}}{\varepsilon_{i}^{l}},$
where the inequality follows from the induction hypothesis (47). For the second term on the right-hand side of (48) a straightforward manipulation (similar to the base case) leads us to
(50) $\left|K_{n}\hat{\pi}_{n-1}^{l}(f)-\hat{K}_{n}^{l}\hat{\pi}_{n-1}^{l}(f)\right|$ $\leq$ $\|f\|_{\infty}\left[\int_{{\mathcal{C}}_{n}^{l}}\int_{{\mathcal{C}}_{n-1}^{l}}K_{n}(x^{\prime},{\sf d}x)\frac{1-\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)}{\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}\bar{x})}\hat{\pi}_{n-1}^{l}({\sf d}x^{\prime})\right$.
$+\left.\int_{\overline{{\mathcal{C}}_{n}^{l}}}\int_{{\mathcal{C}}_{n-1}^{l}}K_{n}(x^{\prime},{\sf d}x)\hat{\pi}_{n-1}^{l}({\sf d}x^{\prime})\right]$
and the inequality (50) together with Assumption 3.3 readily yields
(51) $\left|K_{n}\pi_{n}^{l}(f)-\hat{K}_{n}^{l}\hat{\pi}_{n-1}^{l}(f)\right|\leq\|f\|_{\infty}\left[\frac{1-\varepsilon_{n}^{l}}{\varepsilon_{n}^{l}}+1-\varepsilon_{n}^{l}\right]\leq 2\|f\|_{\infty}\frac{1-\varepsilon_{n}^{l}}{\varepsilon_{n}^{l}},$
where we have used the fact that $\int_{{\mathcal{C}}_{n}^{l}}\int_{{\mathcal{C}}_{n-1}^{l}}K_{n}(x^{\prime},{\sf d}x)\hat{\pi}_{n-1}^{l}({\sf d}x^{\prime})\leq 1$ and $\int_{\overline{{\mathcal{C}}_{n}^{l}}}\int_{{\mathcal{C}}_{n-1}^{l}}K_{n}(x^{\prime},{\sf d}x)\hat{\pi}_{n-1}^{l}({\sf d}x^{\prime})=1-\int_{{\mathcal{C}}_{n}^{l}}\int_{{\mathcal{C}}_{n-1}^{l}}K_{n}(x^{\prime},{\sf d}x)\hat{\pi}_{n-1}^{l}({\sf d}x^{\prime})$. Taking together (49) and (51), we arrive at
(52) $\left|\xi_{n}(f)-\hat{\xi}_{n}^{l}(f)\right|\leq\sum_{i=0}^{n}\bar{c}_{i}\frac{1-\varepsilon_{i}^{l}}{\varepsilon_{i}^{l}},$
where $\bar{c}_{i}=\|f\|_{\infty}\tilde{c}_{i}$, for $i<n$, and $\bar{c}_{n}=2\|f\|_{\infty}$.
To complete the proof, we note that Bayes’ rule yields
(53) $\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)$ $=$ $\frac{\xi_{n}(g_{n}f)}{\xi_{n}(g_{n})}-\frac{\hat{\xi}_{n}^{l}(g_{n}f)}{\hat{\xi}_{n}^{l}(g_{n})}\pm\frac{\hat{\xi}_{n}^{l}(g_{n}f)}{\xi_{n}(g_{n})}$
$=$ $\frac{\xi_{n}(g_{n}f)-\hat{\xi}_{n}^{l}(g_{n}f)}{\xi_{n}(g_{n})}+\hat{\pi}_{n}^{l}(f)\frac{\hat{\xi}_{n}^{l}(g_{n})-\xi_{n}(g_{n})}{\xi_{n}(g_{n})},$
which can be combined with the bound in (52) to obtain the inequality
$\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\leq\sum_{i=0}^{n}c_{i}\frac{1-\varepsilon_{i}^{l}}{\varepsilon_{i}^{l}},\text{~~where~~}c_{i}=\frac{2\|f\|_{\infty}\|g_{n}\|_{\infty}\bar{c}_{i}}{\xi_{n}(g_{n})}.$
We note that, by assumption, $g_{n}>0$ and $\|g_{n}\|_{\infty}$, hence $c_{i}<\infty$ for every $i=0,\ldots,n$. $\Box$
Appendix C Proof of Theorem 3.5
Let $\Phi_{n}$ be the PU operator associated to model $\mathcal{S}$} and let $\Phi_{i:j}=\Phi_{j}\circ\cdots\circ\Phi_{i+1}$, with $\Phi_{i:i}(\pi)=\pi$ for any measure $\pi\in{\mathcal{P}}({\mathcal{X}})$. With this notation, the approximation error can be written as
$\hat{\pi}_{n}^{l}(f)-\pi_{n}(f)=\sum_{i=0}^{n-1}\Phi_{n-i:n}(\hat{\pi}_{n-i}^{l})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l}))(f)+\Phi_{0:n}(\hat{\pi}_{0}^{l})(f)-\Phi_{0:n}(\pi_{0})(f),$
which readily yields the upper bounds [^2] Note that for any probability measures $\alpha$ and $\beta$ one has $\left\|\alpha-\beta\right\|_{\rm tv}=\sup_{|f|\leq 1}\frac{1}{2}|\alpha(f)-\beta(f)|$.
(54) $\sup_{|f|\leq 1}\frac{1}{2}|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|$ $\leq$ $\sum_{i=0}^{n-1}\left\|\Phi_{n-i:n}(\hat{\pi}_{n-i}^{l})-\Phi_{n-i:n}(\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l}))\right\|_{\rm tv}$
$+\left\|\Phi_{0:n}(\hat{\pi}_{0}^{l})-\Phi_{0:n}(\pi_{0})\right\|_{\rm tv},$
$\leq$ $\sum_{i=0}^{n-1}r_{i}\left\|\hat{\pi}_{n-i}^{l}-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})\right\|_{\rm tv}+r_{n}\left\|\pi_{0}-\hat{\pi}_{0}^{l}\right\|_{\rm tv}$
$=$ $\frac{1}{2}\left[\sum_{i=0}^{n-1}r_{i}\sup_{|f|\leq 1}|\hat{\pi}_{n-i}^{l}(f)-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})(f)|+r_{n}|\pi_{0}(f)-\hat{\pi}_{0}^{l}(f)|\right],$
where the second inequality follows from the stability Assumption 3.3.
For the last term on the right-hand side of (54), we obtain
(55) $\sup_{|f|\leq 1}|\pi_{0}(f)-\hat{\pi}_{0}^{l}(f)|\leq 2\frac{1-\varepsilon^{l}}{\varepsilon^{l}}$
using Assumption 3.3 and the same argument that leads to expression (46). As for the $i$-th term in the sum of (54), applying the Bayes’ rule in the same way as in Eq. (53) yields
(56) $\sup_{|f|\leq 1}|\hat{\pi}_{n-i}^{l}(f)-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})(f)|$ $=$ $\sup_{|f|\leq 1}\left|\frac{\hat{K}_{n-i}^{l}\hat{\pi}_{n-i-1}^{l}(g_{n-i}f)-K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i}f)}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}\right$.
$\left.+\hat{\pi}_{n-i}^{l}(f)\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})-\hat{K}_{n-i}^{l}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}\right|$
$\leq$ $\frac{2\int_{{\mathcal{X}}}\int_{{\mathcal{C}}_{n-i-1}^{l}}g_{n-i}(x)|K_{n-i}(x^{\prime},{\sf d}x)-\hat{K}_{n-i}^{l}(x^{\prime},{\sf d}x)|\hat{\pi}_{n-i-1}^{l}({\sf d}x^{\prime})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}.$
In (56), the absolute difference between the kernels is
(57) $|K_{n-i}(x^{\prime},{\sf d}x)-\hat{K}_{n-i}^{l}(x^{\prime},{\sf d}x)|=K_{n-i}(x^{\prime},{\sf d}x)\quad\text{for}\quad{\sf d}x\subset\overline{{\mathcal{C}}_{n-i}^{l}},$
and
(58) $|K_{n-i}(x^{\prime},{\sf d}x)-\hat{K}_{n-i}^{l}(x^{\prime},{\sf d}x)|=K_{n-1}(x^{\prime},{\sf d}x)\frac{1-\int_{{\mathcal{C}}_{n-i}^{l}}K_{n-i}(x^{\prime},{\sf d}\bar{x})}{\int_{{\mathcal{C}}_{n-i}^{l}}K_{n-i}(x^{\prime},{\sf d}\bar{x})}\leq K_{n-1}(x^{\prime},{\sf d}x)\frac{1-\varepsilon^{l}}{\varepsilon^{l}}$
for ${\sf d}x\subset{\mathcal{C}}_{n-i}^{l}$, with the inequality in (57) following form Assumption 3.3. Hence, splitting the the first integral on the right-hand side of (56) over the partition ${\mathcal{X}}={\mathcal{C}}_{n-i}^{l}\cup\overline{{\mathcal{C}}_{n-i}^{l}}$ and combining Eq. (57) and inequality (58) we arrive at
(59) $\sup_{|f|\leq 1}|\hat{\pi}_{n-i}^{l}(f)-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})(f)|$ $\leq$ $\frac{1-\varepsilon^{l}}{\varepsilon^{l}}\times\frac{\int_{{\mathcal{C}}_{n-i}^{l}}\int_{{\mathcal{C}}_{n-i-1}^{l}}g_{n-i}(x)K_{n-1}(x^{\prime},{\sf d}x)\hat{\pi}_{n-i-1}^{l}({\sf d}x^{\prime})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}$
$+\frac{\int_{\overline{{\mathcal{C}}_{n-i}^{l}}}\int_{{\mathcal{C}}_{n-i-1}^{l}}g_{n-i}(x)K_{n-1}(x^{\prime},{\sf d}x)\hat{\pi}_{n-i-1}^{l}({\sf d}x^{\prime})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}$
$=$ $\frac{1-\varepsilon^{l}}{\varepsilon^{l}}\times\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{{\mathcal{C}}_{n-i}^{l}}g_{n-i})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}+\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{\overline{{\mathcal{C}}_{n-i}^{l}}}g_{n-i})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}.$
However, by assumption, $\inf_{i}g_{i}\geq\zeta>0$ and $\sup_{i}\|g_{i}\|_{\infty}\leq\|g\|_{\infty}<\infty$, hence
(60) $\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{{\mathcal{C}}_{n-i}^{l}}g_{n-i})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}\leq\frac{\|g\|_{\infty}}{\zeta}<\infty,$
while
(61) $\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{\overline{{\mathcal{C}}_{n-i}^{l}}}g_{n-i})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}\leq\frac{\|g\|_{\infty}}{\zeta}K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{\overline{{\mathcal{C}}_{n-i}^{l}}})\leq\frac{\|g\|_{\infty}}{\zeta}\times\frac{1-\varepsilon^{l}}{\varepsilon^{l}},$
where the second inequality in (61) follows from Assumption 3.3 (and the fact that $\varepsilon^{l}\leq 1$).
Substituting (60) and (61) back into (59) yields
(62) $\sup_{|f|\leq 1}|\hat{\pi}_{n-i}^{l}(f)-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})(f)|\leq\frac{1-\varepsilon^{l}}{\varepsilon^{l}}\times\frac{2\|g\|_{\infty}}{\zeta}$
and using the inequalities (55) and (62) in expression (54) we obtain
$\sup_{|f|\leq 1}|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|\leq 2\frac{\|g\|_{\infty}}{\zeta}\times\frac{1-\varepsilon^{l}}{\varepsilon^{l}}\times\sum_{i=0}^{n}r_{i}.$
Since $\sum_{i=0}^{\infty}r_{i}<\infty$ by Assumption 3.3, the inequality (15) in the statement of Theorem 3.5 holds with constant $c=2\frac{\|g\|_{\infty}}{\zeta}\sum_{i=0}^{\infty}r_{i}<\infty$ independent of $n$ and $l$. $\Box$
Appendix D Proof of Theorem 3.6
First, we note that the subsets ${\mathcal{C}}_{n}^{l}\subseteq{\mathcal{X}}$ in the constraint ${\mathcal{C}}^{l}$ have been defined to be compact. If every likelihood $g_{n}$ is continuous, with $\lim_{\|x\|\to\infty}g_{n}(x)=0$, then its superlevel sets are compact, and hence it is possible to choose the sets ${\mathcal{C}}_{n}^{l}$ in the constraint ${\mathcal{C}}^{l}$ as superlevel sets of $g_{n}$.
We follow the same argument as in the proof of Theorem 3.5 up to inequality (59). Then we note that
(63) $\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{\overline{{\mathcal{C}}_{n-i}^{l}}}g_{n-i})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})}\leq\frac{\sup_{x\in\overline{{\mathcal{C}}_{n-i}^{l}}}g_{n-i}(x)}{\inf_{x\in{\mathcal{C}}_{n-i}^{l}}g_{n-i}(x)}\times\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{\overline{{\mathcal{C}}_{n-i}^{l}}})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{{\mathcal{C}}_{n-i}^{l}})},$
and, moreover,
(64) $\frac{\sup_{x\in\overline{{\mathcal{C}}_{n-i}^{l}}}g_{n-i}(x)}{\inf_{x\in{\mathcal{C}}_{n-i}^{l}}g_{n-i}(x)}\leq 1\quad\text{and}\quad\frac{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{\overline{{\mathcal{C}}_{n-i}^{l}}})}{K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{{\mathcal{C}}_{n-i}^{l}})}\leq\frac{1-\varepsilon^{l}}{\varepsilon^{l}}$
because ${\mathcal{C}}_{n-i}^{l}$ is a superlevel set of $g_{n-i}$ and Assumption 3.3, respectively. Substituting (64) into (63) and the result into (59) yields
(65) $\sup_{|f|\leq 1}|\hat{\pi}_{n-i}^{l}(f)-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})(f)|\leq 2\frac{1-\varepsilon^{l}}{\varepsilon^{l}},$
where we have also used the fact that $K_{n-i}\hat{\pi}_{n-i-1}^{l}(\mathbbm{1}_{{\mathcal{C}}_{n-i}^{l}}g_{n-i})\leq K_{n-i}\hat{\pi}_{n-i-1}^{l}(g_{n-i})$. Combining (65), (55) and (54) we arrive at
$\sup_{|f|\leq 1}|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|\leq 2\frac{1-\varepsilon^{l}}{\varepsilon^{l}}\sum_{i=0}^{n}r_{i}.$
Since $\sum_{i=0}^{\infty}r_{i}<\infty$ by Assumption 3.3, inequality (16) in the statement of Theorem 3.6 holds with constant $c=2\sum_{i=0}^{\infty}r_{i}<\infty$ independent of $n$ and $l$.
Appendix E Proof of Theorem 3.7
The argument is exactly the same as in the proof of Theorem 3.6, except that Assumption 3.3 yields the time-dependent bounds
(66) $\sup_{|f|\leq 1}|\pi_{0}(f)-\hat{\pi}_{0}^{l}(f)|\leq 2\frac{1-\varepsilon_{0}^{l}}{\varepsilon_{0}^{l}}\quad\text{and}\quad\sup_{|f|\leq 1}|\hat{\pi}_{n-i}^{l}(f)-\Phi_{n-i}(\hat{\pi}_{n-i-1}^{l})(f)|\leq 2\frac{1-\varepsilon_{n-i}^{l}}{\varepsilon_{n-i}^{l}}.$
Substituting (66) back into (54) yields
$\sup_{|f|\leq 1}|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|\leq 2\sum_{i=0}^{n}\frac{1-\varepsilon_{n-i}^{l}}{\varepsilon_{n-i}^{l}}r_{i}.\quad\Box$
Appendix F Proof of Theorem 3.9
F.1 An auxiliary lemma
Let $\pi_{n}^{N}({\sf d}x)=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{n}^{i}}({\sf d}x)$ be the approximation of $\pi_{n}$ constructed by way of the standard particle filter.
Lemma F.1.
Let Assumptions 3.2 and 3.3 hold. If, in addition, there is some $\zeta>0$ such that $\inf_{n\geq 1}g_{n}>\zeta$, then, for any $p\geq 1$,
$\sup_{n\geq 1,|f|\leq 1}\|\pi_{n}(f)-\pi_{n}^{N}(f)\|_{p}\leq\frac{\bar{c}}{\sqrt{N}}$
for some constant $\bar{c}<\infty$ independent of $N$.
Proof F.2.
Let $f:{\mathcal{X}}\mapsto\mathbb{R}$ be a measurable test function, with $\|f\|_{\infty}<\infty$. The difference $\pi_{n}(f)-\pi_{n}^{N}(f)$ can be expanded as
$\pi_{n}^{N}(f)-\pi_{n}(f)$ $=$ $\sum_{i=0}^{n-1}\Phi_{n-i:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n-i:n}\left(\Phi_{n-i}\left(\pi_{n-i-1}^{N}\right)\right)(f)$
$+\Phi_{0:n}\left(\pi_{0}^{N}\right)(f)-\Phi_{0:n}\left(\pi_{0}\right)(f).$
and a straightforward application of Minkowski’s inequality yields
(67) $\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}$ $\leq$ $\sum_{i=0}^{n-1}\left\|\Phi_{n-i:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n-i:n}\left(\Phi_{n-i}\left(\pi_{n-i-1}^{N}\right)\right)(f)\right\|_{p}$
$+\left\|\Phi_{0:n}\left(\pi_{0}^{N}\right)(f)-\Phi_{0:n}\left(\pi_{0}\right)(f)\right\|_{p}.$
for any $p\geq 1$.
For $i=0$ in (67), we have
(68) $\left\|\Phi_{n:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n:n}\left(\Phi_{n}\left(\pi_{n-1}^{N}\right)\right)(f)\right\|_{p}=\|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\|_{p}$
and it is straightforward to show that
$\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)$ $=$ $\frac{\xi_{n}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n}f)}-\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n}f)}\pm\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n}f)}$
$=$ $\frac{\xi_{n}^{N}(g_{n}f)-K_{n}\pi_{n-1}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}+\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\frac{K_{n}\pi_{n-1}^{N}(g_{n})-\xi_{n}^{N}(g_{n})}{\xi_{n}^{N}(g_{n})}.$
Since $\left|\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\right|=|\Phi_{n}(\pi_{n-1}^{N})(f)|\leq\|f\|_{\infty}$, we readily see that
$\left|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\right|\leq\frac{2\|f\|_{\infty}}{\xi_{n}^{N}(g_{n})}\left|\xi_{n}^{N}(g_{n})-K_{n}\pi_{n-1}^{N}(g_{n})\right|.$
From assumption $\inf_{n}g_{n}>\zeta$, we obtain $\xi_{n}^{N}(g_{n})\geq\zeta>0$ and a straightforward application of the Marcinkiewicz-Zygmund inequality [^3] Note that $K_{n}\pi_{n-1}^{N}({\sf d}x_{n})=\frac{1}{N}\sum_{j=1}^{N}K_{n}(x_{n-1}^{j},{\sf d}x_{n})$ and $\xi_{n}^{N}({\sf d}x_{n})=\frac{1}{N}\sum_{j=1}^{N}\delta_{\bar{x}_{n}^{j}}({\sf d}x_{n})$, where $\bar{x}_{n}^{j}\sim K_{n}(x_{n-1}^{j},{\sf d}x_{n})$. If we let $\bar{\mathcal{F}}_{n-1}$ be the $\sigma$-algebra generated by $\{x_{0:n-1}^{1:N},\bar{x}_{1:n-1}^{1:N}\}$, it is clear that $\mathbb{E}[f(\bar{x}_{n}^{j})|\bar{\mathcal{F}}_{n}]=\int f(x_{n})K_{n}(x_{n-1}^{j},{\sf d}x_{n})$ for every j=1,,Nj=1,...,N. yields
(69) $\left\|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\right\|_{p}\leq\frac{2c_{p}\|f\|_{\infty}\|g\|_{\infty}}{\zeta\sqrt{N}},$
where $c_{p}$ is a constant that depends only on $p$ (but is independent of $\xi_{n}^{N}$ and, in particular, of $N$) and we have used the bound $|g_{n}|\leq\|g\|_{\infty}$ from Assumption 3.2.
For $1\leq i\leq n-1$ in (67), let us first note that Assumption 3.3 leads to
(70) $\left\|\Phi_{n-i:n}(\pi_{n-i}^{N})-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))\right\|_{\rm tv}$ $\leq$ $r_{i}\left\|\Phi_{n-i+1}(\pi_{n-i}^{N})-\Phi_{n-i:n-i+1}(\pi_{n-i-1}^{N})\right\|_{\rm tv}$
$=$ $r_{i}\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{n-i+1}(\pi_{n-i}^{N})(f)\right$.
$\quad\quad\quad\quad\quad\left.-\Phi_{n-i+1}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f).\right|$
However, for any $j\geq 2$, one can show that
(71) $\Phi_{j}(\pi_{j-1}^{N})(f)$ $=$ $\frac{\int f(x_{j})g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})\xi_{j-1}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})\xi_{j-1}^{N}({\sf d}x_{j-1})}$
$=$ $\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
and, similarly,
(72) $\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)$ $=$ $\frac{\int f(x_{j})g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}$
$=$ $\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)},$
where $K_{j}(g_{j}f)(x_{j-1})=\int f(x_{j})g_{j}(x_{j})K_{j}(x_{j-1},{\sf d}x_{j})$. Combining (71) and (72) yields
$\Phi_{j}(\pi_{j-1}^{N})(f)-\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)$ $=$
$\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}-\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}\pm\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$ $=$
$\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)-K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
(73) $+\rho_{j}(f)\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)-\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
where
$\rho_{j}({\sf d}x_{j})=\frac{g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}$
is a probability measure. From (73) and the fact that $|\rho_{j}(f)|\leq\|f\|_{\infty}$ we readily find the inequality
(74) $\left|\Phi_{j}(\pi_{j-1}^{N})(f)-\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)\right|\leq\frac{2\|f\|_{\infty}\left|\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)-K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)\right|}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
and, combining (70) and (74) with j=ni+1j=n-i+1 yields
$\sup_{|f|\leq 1}\left|\Phi_{n-i:n}(\pi_{n-i}^{N})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f)\right|$ $\leq$
(75) $2r_{i}\frac{\left|\xi_{n-i}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)-K_{n-i}\pi_{n-i-1}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)\right|}{\xi_{n-i}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)}.$
Recalling, again, the uniform bounds on the likelihoods, $\zeta<g_{n}<\|g\|_{\infty}$, we realise that
(76) $\zeta^{2}<K_{n-i+1}(g_{n-i+1})g_{n-i}<\|g\|^{2}.$
Then, combining (75) with the Marcinkiewicz-Zygmund inequality, applied in the same way as in (69), we readily arrive at the $L^{p}$ bound
(77) $\left\|\Phi_{n-i:n}(\pi_{n-i}^{N})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f)\right\|_{p}\leq\frac{2r_{i}c_{p}\|g\|_{\infty}^{2}}{\zeta^{2}\sqrt{N}}.$
that holds for any real measurable test function $f:{\mathcal{X}}\mapsto\mathbb{R}$ with $\|f\|_{\infty}\leq 1$.
The calculation for the last term on the right hand side of (67) is very similar. Assumption 3.3 yields
(78) $\sup_{|f|\leq 1}\frac{1}{2}\left|\Phi_{0:n}(\pi_{0}^{N})(f)-\Phi_{0:n}(\pi_{0})(f)\right|\leq r_{n}\sup_{|f|\leq 1}\frac{1}{2}\left|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right|$
and
$\left|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right|$ $\leq$
$\left|\frac{K_{1}\pi_{0}^{N}(g_{1}f)-K_{1}\pi_{0}(g_{1}f)}{K_{1}\pi_{0}(g_{1})}+K_{1}\pi_{0}(f)\frac{K_{1}\pi_{0}(g_{1})-K_{1}\pi_{0}^{N}(g_{1})}{K_{1}\pi_{0}(g_{1})}\right|$ $\leq$
(79) $\frac{2\|f\|_{\infty}}{K_{1}\pi_{0}(g_{1})}\left|K_{1}\pi_{0}^{N}(g_{1})-K_{1}\pi_{0}(g_{1})\right|.$
Using again the bounds $\zeta<g_{1}<\|g\|_{\infty}$ and the Marcinkiewicz-Zygmund inequality, we arrive at
(80) $\left\|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right\|_{p}\leq\frac{2\|f\|_{\infty}\|g\|_{\infty}c_{p}}{\zeta\sqrt{N}},$
where the constant $c_{p}<\infty$ depends only on $p\geq 1$. Hence, for any measurable and real test function $f$ with $\|f\|_{\infty}\leq 1$, the combination of (78), (79) and (80) yields
(81) $\left\|\Phi_{0:n}(\pi_{0}^{N})(f)-\Phi_{0:n}(\pi_{0})(f)\right\|_{p}\leq\frac{4r_{n}c_{p}\|g\|_{\infty}}{\zeta\sqrt{N}}.$
Finally, recalling Assumption 3.3 (in particular, $r_{0}=1$ and $\sum_{i=0}^{\infty}r_{i}<\infty$) and taking together (67), (69), (77) and (81) we arrive at
$\sup_{|f|\leq 1}\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}\leq\frac{4c_{p}\|g\|_{\infty}^{2}\sum_{i=0}^{\infty}r_{i}}{\zeta^{2}}\times\frac{1}{\sqrt{N}}.$
F.2 Proof of Theorem 3.9
For the first part of the theorem, let us take the triangle inequality
(82) $\|\pi_{n}(f)-\pi_{n}^{l,N}(f)\|_{p}\leq|\pi_{n}(f)-\pi_{n}^{l}(f)|+\|\pi_{n}^{l}(f)-\pi_{n}^{l,N}(f)\|_{p}.$
Assumptions 3.3 and 3.3, and the uniform bounds $\zeta<g_{n}<\|g\|_{\infty}$ on the likelihoods yield, by way of Theorem 3.5, the first term on the r.h.s. of (18). In particular, there is some constant $c<\infty$ such that
(83) $\sup_{|f|\leq 1}|\pi_{n}(f)-\pi_{n}^{l}(f)|\leq c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}.$
As for the term $\|\pi_{n}^{l}(f)-\pi_{n}^{l,N}(f)\|_{p}$ on the right-hand side of (82), we note that Assumption 3.2 and (17) enable us to use Theorem 3.1 which, in turn, ensures that the PU operator $\hat{\Phi}_{n}^{l}$ is mixing with some constant $\gamma_{l}>0$. Then, Theorem 2.6 yields
(84) $\sup_{|f|\leq 1,n\geq 0}\|\pi_{n}^{l}(f)-\pi_{n}^{l,N}(f)\|_{p}\leq\frac{c_{p}\|g\|_{\infty}}{\zeta\gamma_{l}^{6}\sqrt{N}}$
for some $c_{p}<\infty$. Taking (82), (83) and (84) together yields inequality (18) in the statement of Theorem 3.9.
Inequality (19) is straightforward. Simply note that, for fixed $l$, $\gamma_{l}>0$ and then let $N\to\infty$ in (18).
For the third part of the theorem, we iterate the triangle inequality to arrive at
(85) $\lim_{l\to\infty}\|\pi_{n}(f)-\pi_{n}^{l,N}(f)\|_{p}$ $\leq$ $\|\pi_{n}(f)-\pi_{n}^{N}(f)\|_{p}+\lim_{l\to\infty}\|\pi_{n}^{N}(f)-\pi_{n}^{l}(f)\|_{p}$
$+\|\pi_{n}^{l}(f)-\pi_{n}^{l,N}(f)\|_{p},$
and then we note that $\sup_{l\geq 1}|\pi_{n}^{l}(f)|\vee|\pi_{n}^{l,N}(f)|\leq\|f\|_{\infty}<\infty$, hence Lebesgue’s dominated convergence theorem yields
(86) $\lim_{l\to\infty}\|\pi_{n}^{N}(f)-\pi_{n}^{l}(f)\|_{p}=\|\pi_{n}^{N}(f)-\lim_{l\to\infty}\pi_{n}^{l}(f)\|_{p}=\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}$
and, similarly,
(87) $\lim_{l\to\infty}\|\pi_{n}^{l}(f)-\pi_{n}^{l,N}(f)\|_{p}=\left\|\lim_{l\to\infty}\pi_{n}^{l}(f)-\pi_{n}^{l,N}(f)\right\|_{p}=\|\pi_{n}(f)-\pi_{n}^{N}(f)\|_{p}.$
Taking together (85), (86) and (87) we obtain
$\lim_{l\to\infty}\|\pi_{n}(f)-\pi_{n}^{l,N}(f)\|_{p}\leq 3\|\pi_{n}(f)-\pi_{n}^{N}(f)\|_{p}$
and Lemma F.1 ensures the existence of some constant $\bar{c}<\infty$ such that
$\sup_{|f|\leq 1,n\geq 0}\|\pi_{n}(f)-\pi_{n}^{N}(f)\|\leq\frac{\bar{c}}{\sqrt{N}}.$
Hence, (20) is satisfied.  $\Box$